### v2

# Mixture of Cluster-conditional LoRA Experts for Vision-language Instruction Tuning

Abstract
Instruction tuning of the Large Vision-language Models (LVLMs) has revolutionized the development of versatile models with zero-shot generalization across a wide range of downstream vision-language tasks. However, diversity of training tasks of different sources and formats would lead to inevitable task conflicts, where different tasks conflicts for the same set of model parameters, resulting in sub-optimal instruction-following abilities. To address that, we propose the Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel Mixture of Experts (MoE) architecture designed to activate the task-customized model parameters based on the instruction clusters. A separate universal expert is further incorporated to improve the generalization capabilities of MoCLE for novel instructions. Extensive experiments on 10 zero-shot tasks demonstrate the effectiveness of MoCLE.
$*{}^{*}$ Equal contribution. ${}^{\dagger}$ Corresponding author.
1 Introduction
There has been a continuously increasing trend to develop intelligent assistants that can follow human instructions [4, 42, 8], and instruction tuning [55] has become a successful paradigm leveraging large-scale well-formatted instruction data to empower Large Language Models (LLMs) to follow human instructions, demonstrating strong generalization for novel unseen tasks [36]. Similarly, efforts have been made to introduce similar capabilities to Large Vision-language Models (LVLMs), including LLaVA [33], MiniGPT-4 [67] and InstructBLIP [11], which connects a trainable visual encoder with a frozen LLM and endows the LVLMs with the instruction-following ability by fine-tuning on multi-modal instruction datasets.
The ability of generalizing to novel unseen instructions necessitates multi-task instruction tuning, i.e., training on a diverse collection of instruction-following tasks, which has been observed in both the LLMs [45, 53, 10] and LVLMs [3, 63, 28]. However, complexity of various instruction tasks brings difficulties for the model fine-tuning. Specifically, Wei et al. [55] observe that under certain model sizes, multi-task instruction tuning even fails to bring performance gains for zero-shot tasks compared to the original models, mainly attributed to the negative transfer phenomenon [62] during multi-task instruction tuning, where the model struggles to optimize the losses of multiple conflicted tasks, leading to sub-optimal performance for these tasks.
Figure 1: Performance of the instruction-finetuned LVLMs on zero-shot tasks, where larger values indicate better performance. Only 3 out of 8 tasks benefit from instruction tuning from all the data, while the task experts show better performance on the other 5 tasks (i.e., Flickr 30K, HM, ScienceQA, IconQA, and VizWiz).
Similarly, tasks for vision-language instruction tuning (e.g., visual question answering and image captioning) focus on different perspectives of LVLMs, and thus, resulting in conflicts when sharing all the parameters as most studies adopt. In our preliminary studies shown in Figure 1, we first split the instruction data into two disjoint subsets, including “cap” for image captioning and “vqa” for visual question answering. We then train InstructBLIP [11] using LoRA [18] on 3 kinds of data: “cap”, “vqa” and “full” (the full data) separately, to obtain 3 set of parameters (i.e., task experts). Following the held-out evaluation protocol in [11], we evaluate these experts on the unseen datasets/tasks with the best expert. As shown in Figure 1, on 5 out of the 8 downstream tasks, InstructBLIP instruction-tuned on all the data is outperformed by the task expert that is finetuned on only a subset of data. Among the 5 tasks, Flickr30k belongs to “cap”, and SciQA, IconQA and VizWiz belong to “vqa”, revealing that instruction tuning on the similar tasks brings positive transfer to downstream tasks, while training on the full data with dissimilar tasks could hurt generalization performance.
The task experts above is a naive solution to negative transfer, where we manually partition the training tasks and train each expert individually. However, there are several limitations of this strategy, (1) The taxonomies such as “vqa” and “cap” require human expertise and become difficult to scale as the number of tasks grows; (2) The ability to generalize to unseen tasks is inhibited, as we do not know which expert to choose for novel tasks and some new tasks might benefit from multiple training tasks (e.g., VSR and TextVQA as in Figure 1). In this regard, specialization and generalization of LMMs becomes a dilemma.
Our work aims to develop an automatic and practical partition strategy as well as a network architecture that can strike a balance between specialization and generalization. We propose Mixture of Cluster-conditional LoRA Experts (MoCLE) for vision-language instruction tuning. Under the proposed MoCLE framework, we first cluster instructions of all the training data into several centroids via a pretrained clustering model. In this way, similar tasks that could bring positive transfer to each other are grouped together in a cluster (more justification for the use of instruction clusters is detailed in Sec. 3.3). Then using the cluster as a condition, at each layer of the model, a router dispatches the input data to one of the specialized task experts and also a universal expert that is shared among all the data. Finally, these two set of outputs are combined and passed to the next layer. As we activate a specialized expert for a group of similar tasks, tasks that are less alike are learned via separate experts, mitigating task conflicts. Meanwhile, since the universal expert trained on all tasks also contributes to the model outputs, we can enjoy generalization and specialization simultaneously.
We validate the effectiveness of MoCLE on the state-of-the-art InstructBLIP [11]. MoCLE achieves remarkable improvement on the held-out tasks [11] including image captioning (e.g., Flickr30K [58]), text-based visual question answering (e.g., TextVQA [50]), knowledge-based visual question answering (e.g., ScienceQA [38]) and novel visual reasoning (e.g., Visual Spatial Reasoning [32]).
The main contributions of this work contain three parts.
  1. 1.
    We identify the negative transfer phenomenon [34, 65] of task conflicts during instruction tuning of LVLMs.
  2. 2.
    We propose the Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel parameter-efficient finetuning framework suitable for the vision-language instruction tuning, to mitigate task conflicts and enjoy the benefits of huge data training simultaneously.
  3. 3.
    The proposed MoCLE method achieves remarkable performance gains on the unseen held-out tasks compared with a strong baseline model InstructBLIP [11].
2 Related Work
Multi-task Instruction Tuning.
Instruction tuning, as elucidated in [45, 55], is a technique where a language model is finetuned over a collection of tasks and instruction templates that tells the model the intention of the task. The primary goal of this technique is to teach the model to understand the relationship between the instruction and the input/output pairs, so that the model could generalize to unseen task given new instructions. Increasing the number of diverse instructions [45], the number of tasks [53, 10], and diversity of data [66] have been shown to improve the performance. However, Wei et al. [55] observed under certain model sizes, instruction tuning of language models fail to bring performance gains to unseen tasks when compared with untuned base models. They argue this is due to the fact all model capacity is used to learn the mixture of instruction tuning tasks. In this work, we aim to mitigate such issue in vision-language instruction tuning by specialized experts.
Large Multi-modal Models.
To inherit strong reasoning and generalization abilities of LLMs and meanwhile understand visual inputs, recent works including Flamingo [2], BLIP-2 [26], LLaMA-adapter [61] connect LLMs to the visual modality with a frozen visual encoder and a trainable module, which is then finetuned with plain image-text pairs. To further improve instruction following abilities and response quality, LLaVA finetunes the whole LLM on vision-language conversational data generated by GPT-4. MiniGPT-4 performs training on ChatGPT-generated image captions instead of traditional caption data that are much shorter. InstructBLIP collects a wide range of academic VQA datasets and manually crafts 10 to 15 distinct instruction templates for each kind of task. In this paper, the proposed MoCLE is built on InstructBLIP as the base model.
Figure 2: Overall pipeline of MoCLE. The left part is to cluster training data (i.e., a collection of vision-language tasks with diverse instructions) into several groups based on their instructions via $k$-means clustering algorithm. On the right is the MoCLE framework. $\mathbf{x}$ are the hidden representations of the sample (e.g., a sequence of tokens for images, instructions and the raw text inputs in the LLM). The input $\mathbf{x}$ is handled by 3 modules, including the universal expert, the task expert with the largest gating value, and the linear module of the LLM. After that, the outputs of these modules are aggregated and passed to the next layer. Only experts, the cluster embedding and the gate parameters are updated during training, while parameters of the LLM are frozen. We insert such structure into the query and value projection of self-attention blocks of each layer by default.
Mixture of Experts.
MoE models [20, 21, 47] are considered as an effective way to increasing the model capacity in terms of parameter size. In MoE, certain parts of the model are activated while the computation is kept the same or close to its dense counterpart. Recently, it has been thoroughly investigated in the field of computer vision [44, 65], natural language processing [24, 12] and multi-modal learning [41, 48]. Among those studies, MoCE [65] tackles negative transfer in the self-supervised visual pre-training [6, 7] using a mixture of cluster experts. Different from MoCE that needs to be finetuned on downstream tasks, the proposed MoCLE method studies the zero-shot generalization to unseen tasks. Another line of works combines MoE with adapters. AdaMix [54] randomly selects adapters during training and uses the average parameters of those experts during the inference. Zadouri et al. [60] proposes mixture of LoRAs for instruction tuning of LLMs, achieving comparable results with the fully-finetuned counterpart. However, neither of them explicitly addresses task conflicts and generalization to unseen tasks, which is the focus of our work.
3 Methodology
In this section, we start with the formulation of LVLM instruction tuning and the analysis of the limitations of task experts. We then introduce the proposed MoCLE in detail. Figure 2 shows the overall framework of MoCLE.
3.1 Problem Formulation
Suppose that there are a set of datasets which are divided into held-in and held-out datasets, a large vision-language model is first fine-tuned on the held-in and then evaluated in a zero-shot manner on the held-out datasets. To unify input-output formats and promotes generalization in instruction tuning, several task templates $\{T_{i}\}$ are designed to wrap the raw inputs, i.e., a pair of text and image $(X_{\text{txt}},X_{\text{img}})$ from the corresponding dataset, where $T_{i}$ denotes the $i$-th template. For example, “Given the image, answer the question with no more than three words. {Question}” is a template for the visual question answering tasks. An instruction $I\equiv T_{i}(X_{\text{txt}})$ wraps the text input using the template.
3.2 Limitation of Task Experts
As observed in Figure 1, allocating task experts based on the categories of the training datasets brings positive transfer to similar downstream tasks. However, there are several limitations of this solution. The first limitation is that taxonomies such as “vqa” and “cap” require delicate human designs with expertise in the task performed. The second limitation is that when a new task comes, we do not know which expert to use. Further, with tasks trained in an isolated manner, knowledge sharing among different tasks is prevented, compromising generalization to unseen tasks and instructions as some might benefit from multiple tasks. Hence, we propose the MoCLE in the following sections to alleviate these two limitations.
3.3 Clustering Data by Instructions
The goals of partitioning the training data are two-fold. We target to train a task expert within a collection of similar tasks so as to prevent task conflicts. Meanwhile, we expect new tasks to be assigned to the proper experts based on its cluster without manual intervention. To achieve these, we should find a proper description for the task. Instructions serve as the foundation for LVLMs because they encapsulate the commands and intentions that guide responses of the model. The instructions $I$ in LVLMs consist of templates and the raw inputs. By understanding the templates in the instructions, the model can discern the type of task it needs to perform. The raw inputs further infuse this framework with the substantive content that the model must process and respond to. Therefore, by conducting clustering on the instructions, we can partition training data without human labor into several clusters, each containing a set of similar tasks. We further verify this through experiments later in Sec. 4.5. Moreover, even when novel instruction tasks come, we can obtain a cluster index for them using the pretrained clustering model.
Formally, denoting $\mathcal{E}(\cdot)$ as a pretrained sentence encoder, $\mathbf{e}_{i}=\mathcal{E}(I_{i})$ is the sentence representation of an instruction $I_{i}$, we use the $k$-means clustering algorithm to group all the instructions in the training datasets into $K$ clusters by iteratively minimizing $\sum_{j=1}^{K}\sum_{\mathbf{e}_{i}\in S_{j}}\|\mathbf{e}_{i}-\mathbf{c}_{j}\|^{2}$, where $S_{j}$ denotes the set of instructions assigned to the $j$-th cluster and $\mathbf{c}_{j}$ denotes the centroid of the $j$-th cluster. In each iteration of the $k$-means clustering algorithm, each instruction is assigned to the nearest centroid with all centroids updated to the average of instructions in the corresponding cluster.
3.4 Mixture of Cluster-conditional Experts
In Sec. 1, we defined a static one-to-one mapping between the task categories and expert parameters. For example, the data in the $i$-th cluster are processed by expert-$j$ for any layer at all the time. Such static mapping is sub-optimal since $N$ tasks categories require $N$ task experts, resulting prohibited model sizes as $N$ becomes large. In fact, an expert does not necessarily have to handle only one kind of tasks all the time. That is, at different layers, a task could have different choices of experts. Concretely, consider $L$ layers each with $N$ experts, then there are $N^{L}$ number of choices in total. This would accommodate much more tasks while keep $N$ small. Yet, under such a circumstance, how to choose the expert for a given task remains unclear.
We propose Mixture of Cluster-conditional LoRA Experts (MoCLE) that learns to activate the LoRA expert at each layer given the cluster of the data. Specifically, denoting by $E$ the number of experts, we introduce a gate vector $\mathbf{G}\in\mathbb{R}^{E}$ that, given any input $\mathbf{x}_{i}$, determines the experts to which the input is routed. The gate vector is calculated as,
$\mathbf{G}=\text{top}_{k}\left(\text{softmax}\left(\frac{1}{\tau}\left(\mathbf{W}_{\text{gate}}\mathbf{C}_{[\mathbf{x}_{i}]}+\boldsymbol{\epsilon}\right)\right)\right),$ (1)
where $\text{top}_{k}(\cdot)$ keeps the $k$ largest entries unchanged and sets others to zero and $\mathbf{C}_{[\mathbf{x}]}$, which is shared among all the layers, denotes the learnable embedding of the cluster that $\boldsymbol{x}$ belongs to. This is the key for the model to choose proper task experts for the input data. To endow the clustering embedding with task information, we initialize it to be the centroid of the represented cluster. $\mathbf{W}_{\text{gate}}$ denotes trainable weights of the linear gate, $\boldsymbol{\epsilon}\sim N(0,\frac{1}{E})$ denotes a noise term that is designed to add some randomness to the expert choosing process, which encourages MoCLE to explore multiple combinations of experts during training. $\tau$ denotes a temperature hyperparamter. The output $\mathbf{y}_{i}$ is then computed as a weighted sum of the output of each expert on the input $\mathbf{x}_{i}$ as,
$\mathbf{y}_{i}=\sum_{e=1}^{E}G_{e}\mathbf{W}_{e}\mathbf{x}_{i},$ (2)
where $\mathbf{W}_{e}$ represents the linear projection weights of the $e$-th expert, and $G_{e}$, the $e$-th entry in $\mathbf{G}$, indicates the contribution of the $e$-th expert to $\mathbf{y}_{i}$.
In this formulation, the model learns to choose the expert that is suitable for the data by examining its cluster. In addition, as a task has different choices of expert at different layers, expert number scales well with the number of task clusters. We mitigate task conflicts with only a few experts.
3.5 Universal Expert
However, as shown in Sec. 4.4.1, we find that the formulation in Sec. 3.4 still hurts the generalization ability of the entire model. We hypothesize that this might be attributed to the fact that instruction-tuned models generalize to unseen tasks by training on many instructions, while in our case, each expert sees fewer instructions than the original dense model does. Hence, we propose a universal expert that could learn from all the training data. Specifically, we fix the number of activated experts to 1 (i.e., $k$ equals 1) and define $G_{\text{max}}$ as the maximum element in $\mathbf{G}$. Then the output for all the experts is expressed as,
$\mathbf{y}_{i}=\left(\sum_{e=1}^{E}G_{e}\mathbf{W}_{e}+(1-G_{\text{max}})\mathbf{W}_{u}\right)\mathbf{x}_{i}.$ (3)
In this formulation, we additionally train a universal expert parameterized by $\mathbf{W}_{u}$. Different from task experts that are activated only for specific model inputs, the universal expert is activated for any inputs. The final output is the weighted sum of outputs of one of the experts and the universal expert. The weighting for this universal expert is the complement of that of the chosen task expert. Consequently, the task expert learns distinct skills for certain tasks while the universal expert masters holistic understanding of the training corpus. The synergy between them offers both specialization and generalization for the LVLM.
4 Experiment
In this section, we empirically evaluate MoCLE on various downstream tasks in a zero-shot manner.
4.1 Settings and Implementation Details
In all the experiments, we utilize the InstructBLIP [11] model, particularly its Vicunna-7B variant [64], as the large vision-language model (LVLM) to assess the effectiveness of the proposed method. We encode the instructions of different datasets using the all-MiniLM-L6-v2 variant of the Sentence Transformer model [43] and cluster their embeddings into $64$ clusters via $k$-means clustering algorithm. During the vision-language instruction tuning, we initialize the model based on pre-trained InstructBLIP checkpoints [11] and apply the proposed MoCLE on the LLM (i.e., Vicunna-7B). Specifically, for each transformer layer in the LLM, we insert the MoCLE, which consists of 4 task LoRA experts and one LoRA universal expert, into the query and value projection of self-attention blocks. Inside each expert, we set the rank of the LoRA to 8 and the temperature $\tau$ to $0.05$. During instruction tuning, we only finetune parameters of Q-Former and the LoRA experts while keeping other parts of the LLM frozen.
Flickr 30K GQA VSR IconQA TextVQA Visdial HM VizWiz SciQA image MSVD QA iVQA Trained using private data Flamingo-80B 67.2 - - - 35.0 - 46.4 31.6 - 35.6 40.7 Qwen-VL (7b) 85.8 $*{}^{*}$ $*{}^{*}$ 35.2 67.1 - - Qwen-VL (7b)-chat 81.0 $*{}^{*}$ $*{}^{*}$ 38.9 68.2 - - Trained using grounding data MiniGPT-v2 (7B) - $*{}^{*}$ 60.6 47.7 - - 58.2 32.9 - - - MiniGPT-v2 (7B)-chat - $*{}^{*}$ 62.9 51.5 - - 58.8 53.6 - - - Trained using public data LLaVA (13B) - 41.3 51.2 43.0 - - - - - - - MiniGPT-4 (13B) - 30.8 41.6 37.6 - - - - - - - BLIP-2 (7B) 74.9 38.6 50.0 39.7 40.1 44.9 50.6 25.3 53.8 18.3 27.5 InstructBLIP (7B) 82.4 49.2 54.3 43.1 50.1 45.2 59.6 34.5 60.5 41.8 52.2 Under the same evaluation InstructBLIP (${}^{{\dagger}}$ 82.9 49.2 65.2 43.1 50.9 44.2 64.3 30.3 59.4 41.5 51.7 InstructBLIP (${}^{{\ddagger}}$ 81.3 48.6 60.8 43.4 53.9 42.7 65.3 29.5 62.0 41.4 51.3 InstructBLIP (${}^{{\ddagger}}$+MoCLE 81.9 (+0.6) 49.3(+0.7) 64.7(+3.9) 46.3(+2.9) 57.1(+3.2) 45.3(+2.6) 65.6(+0.3) 30.4(+0.9) 63.9(+1.9) 42.6(+1.2) 53.2(+1.9)
Table 1: Zero-shot results on the held-out datasets. Here, Visdial, HM and SciQA denote Visual Dialog, HatefulMemes and ScienceQA, respectively. For ScienceQA, we only evaluate the set with image context. Following previous works [11], we report the CIDEr score [52] for Flickr30K, accuracy for iVQA, AUC score for HatefulMemes, and Mean Reciprocal Rank (MRR) for Visual Dialog. For all other datasets, we report top-1 accuracy (%). Notice results with * means the training split of the evaluation dataset is used during instruction tuning, thus not zero-shot. Best results obtained by models trained with public data are marked in bold and the second best are underlined. Superscript ${}^{{\dagger}}$ denotes the official InstructBLIP checkpoint. Superscript ${}^{{\ddagger}}$ denotes our reproduced version, into which MoCLE is incorporated.
4.2 Datasets and Instructions
To demonstrate the effectiveness of MoCLE, we follow InstructBLIP to use the same training and evaluation data unless mentioned explicitly. As Web CapFilt [26], one of the training datasets, is not fully available, we can only download a subset of them. Besides this, we utilize all the other training data, including COCO [31], TextCaps [49], VQAv2 [15], OKVQA [39], A-OKVQA [46], OCR-VQA [40], and LLaVA-Instruct150K [33], for instruction tuning.
For evaluation datasets, we follow InstructBLIP but only keep Flickr30K [58], GQA [19], VSR [32], IconQA [37], TextVQA [50], Hateful Memes [22], VizWiz [16], ScienceQA [38], MSVD-QA [56], and iVQA [57]. InstructBLIP originally uses the NoCaps dataset [1] for the image captioning task, which is not adopted here since we cannot reproduce the reported results, probably due to the incomplete Web CapFilt dataset used. In addition, we do not utilize the MSRVTT-QA dataset [56] due to the insufficient computational resources.
The instructions to conduct all the experiments in this paper are from the original InstructBLIP paper. The detailed introductions of the held-in and held-out tasks/datasets are also presented in [11].
4.3 Zero-shot Evaluations
Table 1 presents the experimental results on multiple zero-shot vision-language tasks. We build our method upon the Vicunna-7B variant of InstructBLIP. The InstructBLIP (${}^{{\ddagger}}$ is our reproduced version of InstructBLIP. Compared to this model, the proposed MoCLE method shows considerable performance improvement. Specifically, on conventional tasks such as visual question answering (VQA), we observe a 2.9% improvement on IconQA. On tasks with novel formats such as Visual Spatial Reasoning (VSR), TextVQA and ScienceQA image, we obtain an absolute performance gain of 3.9%, 3.2%, and 1.9%, respectively. This indicates that the proposed MoCLE method facilitates generalization to unseen tasks/instructions since the instructions of VSR, TextVQA and ScienceQA image are unseen during training. In addition, the proposed MoCLE method performs well when compared with existing popular LVLMs including LLaVA (13B) and MiniGPT-4 (13B). For example, on IconQA, the proposed MoCLE method outperforms LLaVA (13B) and MiniGPT-4 (13B) by 3.3% and 8.7%, respectively. In summary, the proposed MoCLE model obtains best results on IconQA, TextVQA, HM, and ScienceQA image across models trained using public data.
4.4 Ablation Studies
In this section, we first ablate the effectiveness of the main components (e.g., task experts and universal expert) in the proposed MoCLE method in Sec. 4.4.1. Then in Sec. 4.4.2 we conduct a thorough analysis to study how the proposed MoCLE method would respond to the change of hyper-parameters (e.g., temperature and the number of clusters and task experts). Notice that for the ablations, we use Flickr30K (Flickr), Hateful Memes (HM), ScienceQA (SciQA), IconQA, Visual Spatial Reasoning (VSR) and TextVQA for evaluation.
Components Flickr HM SciQA IconQA VSR TextVQA
InstructBLIP 81.3 65.3 62.0 43.4 60.8 53.9
+LoRA (rank=8) 81.3 65.1 57.4 44.2 62.8 49.4
+LoRA (rank=64) 81.5 65.2 62.0 43.9 62.6 49.0
+Mixture (Cluster) 81.9 65.4 63.3 46.1 58.9 54.9
+Universal Expert 81.9 65.6 63.9 46.3 64.7 57.1
Table 2: Comparison of individual components of the MoCLE framework in zero-shot vision-language tasks. We include three ablation studies that cumulatively add (i) parameter-efficient finetuning of LLM using LoRA of different ranks, (ii) replacing the single LoRA with its mixture and route inputs by their cluster, and (iii) training with the universal expert. Default settings are marked in gray.
4.4.1 Effects of Different Components
We start from vanilla InstructBLIP (Vicunna-7B) and add the main components in the proposed MoCLE method step by step. As shown in Table 2, naively applying LoRA to InstructBLIP could be problematic. Although we observe minor performance gain in IconQA and VSR, the peformance on ScienceQA and TextVQA are much inferior to InstructBLIP. We speculate that a single LoRA offers very limited flexibility, thus it performs not so well on many tasks. Note that we further conduct an experiment with a rank of 64 (much larger than $r=8$ in the default setting of MoCLE). However, the resulting model is still outperformed by InstructBLIP with less trainable parameters. This indicates that naively increasing the capacity of the network cannot mitigate the task conflicts in the training data. However, when we replace a single LoRA with the Mixture of LoRAs and route the input based on the cluster of their instructions, the resulting model outperforms its predecessor by a large margin on SciQA, IconQA, and TextVQA. Nevertheless, we notice a sharp performance drop in the VSR task. This could be due to the fact that instruction-tuned model generalizes to unseen tasks by training on many instructions, while in our case, each expert sees fewer instructions than the original dense model does. As shown in Table 2, this issue could be alleviated by the use of the universal expert, which is updated by the gradients from all the training data. The same result can be observed in TextVQA, which requires not only the VQA ability but also additional optical character recognition (OCR) skills. Specifically, in the training stage, we have instructions “Question: {}. Short answer:” for all the VQA datasets and “OCR tokens: {}. Write a description for the photo” for TextCaps [49], but the instruction of TextVQA [50] is “OCR tokens: {}, Question: {}. Short answer:”. This requires the model to combine the knowledge from two tasks.
Universal Expert vs. Top-2 Experts.
MoCLE shares similarities with the top-2 variant of MoE. Specifically, they both activate two experts each time. However, the universal expert differs from top-2 MoE by consistently selecting the universal expert, whose contribution is the complement of the top-1 expert. As shown in Table 3, the top-2 MoE model yields inferior results compared to MoCLE. Therefore the naively activating one more expert cannot address the generalization problem. Surprisingly, this variant performs even worse than the MoCLE model without the universal expert reported in Table 2. This might be resulted from task conflicts with more parameters shared.
Flickr HM SciQA IconQA VSR TextVQA
Universal 81.9 65.6 63.9 46.3 64.7 57.1
top-2 82.0 64.7 61.9 45.5 56.3 52.0
Table 3: Ablation study on the universal expert by comparing with either (i) a universal expert that is activated all the time or (ii) expert with the second largest logit, in addition to the top-1 expert.
Gating strategy Flickr HM SciQA IconQA VSR TextVQA
Token 81.7 65.4 61.9 44.0 49.0 46.6
Sentence 82.0 65.1 62.3 45.3 56.6 47.0
Cluster 81.9 65.6 63.9 46.3 64.7 57.1
Table 4: Ablation study on routing inputs based on different conditions. Default settings are marked in gray.
Gating Strategies.
We compare the proposed cluster-conditioned gating strategy with two popular ones: Token-MoLE and Sentence-MoLE, where MoLE denotes the mixture of LoRA experts by applying the MoE technique to LoRA, Token-MoLE obtains the routing decision based on the hidden representations of each token, and Sentence-MoLE averages the representations of the instruction tokens while excluding the visual tokens. As observed in Table 4, Token-MoLE and Sentence-MoLE give inferior results on the evaluation tasks. We speculate this is because that (1) an expert only learns a subset of the tasks, giving poor generalizations and (2) similar tasks are not grouped together by the same expert, resulting in task conflicts within that expert. The later can be evidenced by the routing visualization in Figure 5, where samples in the same dataset are routed to multiple experts instead of a dedicate one.
4.4.2 Hyper-parameter Analysis
In this section, we analyze in details the effect of the hyper-parameter in the MoCLE framework to the performance. Specifically, we investigate the how would MoCLE react to the number of clusters as introduced in Sec. 3.3 and the number of task experts as well as the temperature value in Sec. 3.4. If not otherwise specified, we employ all the components ablated in Sec. 4.4.1 and use a default configuration of 64 clusters with 4 experts and a temperature of 0.05.
# Clusters Flickr HM SciQA IconQA VSR TextVQA
4 81.3 65.2 63.4 45.3 58.3 55.3
16 82.0 64.3 61.7 46.3 67.2 56.9
64 81.9 65.6 63.9 46.3 64.7 57.1
256 82.5 65.6 63.2 46.0 57.0 56.9
Table 5: Ablation study on the number of clusters $K$.
#Task experts Flickr HM SciQA IconQA VSR TextVQA
2 81.4 64.4 62.8 45.6 58.1 50.9
4 81.9 65.6 63.9 46.3 64.7 57.1
8 81.0 65.3 63.5 45.9 65.0 57.4
Table 6: Ablation study on the number of task experts.
Temperature $\tau$ Flickr HM SciQA IconQA VSR TextVQA
0.01 81.6 64.7 61.7 46.3 61.2 55.8
0.05 81.9 65.6 63.9 46.3 64.7 57.1
0.1 82.3 65.7 63.7 45.7 63.0 56.7
0.2 81.9 64.8 62.1 45.2 61.3 54.0
Table 7: Ablation study on the temperature $\tau$ in Eq. (1).
Number of Clusters.
The merit of sentence encoding and clustering lies in reducing the complexity of the instruction, offering robust and semantic-relevant routing process. The number of clusters $K$ in this process could measure such complexity. A extremely small $K$ would diminish the difference of various tasks, result in many different tasks to be processed by the same expert. This could increase the chance of task conflicts. As shown in Table 5, when we cluster the inputs into 4 groups, the resulting model performs poorly on the evaluation tasks. However, as we increase the number of clusters to 16 and 64, we observe considerable performance gains. Nevertheless, too large $K$ introduces unnecessary complexity in the routing process, e.g., a paraphrased instruction gets routed to different experts. So we use 64 clusters as the default setting.
Number of Task Experts.
As demonstrated in Table 6, more task experts usually provides with stronger capacity. Specifically, when only 2 task experts are employed, we observe inferior results on VSR and TextVQA. This model has similar capacity to the single LoRA model in Sec. 4.4.1, where only one LoRA encounters difficulties in fitting a diverse set of tasks. When the number of task experts is increased to 4, the performance gets improved. Specifically, on VSR and TextVQA, the 4-expert variant outperforms the 2-expert counterpart by 6.6% and 6.2%, respectively. Moreover, we also observe 1.2%, 1.1%, and 0.7% performance gains on HM, SciQA and IconQA. When the number of task experts becomes 8, it behaves similarly to the 4-expert case, which indicates that the benefit of increasing capacity converges as we use more task experts. Hence, we use 4 task experts as the default setting.
Figure 3: t-SNE visualization of the instruction encoding. Different colors correspond to different datasets, while the shape of the markers indicates the task category defined manually.
Temperature.
In the proposed MoCLE method, the value of temperature plays an important role in controlling the contribution of the universal expert. Specifically, as shown in Eq. (1), $\tau$ controls the sharpness of the gate distribution, while the output of the universal expert is weighted by $1-G_{\text{max}}$. Therefore, as $\tau$ decreases, $G_{\text{max}}$ increases, and finally the contribution of the universal expert decreases. As shown in Table 7, the results are consistent with our understandings. When a $\tau$ is too small, i.e., 0.01, the model generalizes worse to unseen tasks like SciQA and VSR. Similarly, when $\tau$ is too large, i.e., $0.2$, the role of task experts is diminished, leading to task conflicts. The temperatures of $0.05$ and $0.1$ seem to achieve a balance between specialization and generalization of the model, and we use 0.05 as the default setting for the temperature.
4.5 Visualizations
Figure 4: Clustering assignment of the training datasets when $K=64$. The labels on the vertical axis indicate the names of the datasets and they are arranged such that different datasets of the same task are placed together. Within each labeled block, multiple rows represent different subsets, each of which is associated with a unique task template. The horizontal axis denotes the cluster index to which these subsets are assigned. Darker color indicates more samples of this subset are assigned to certain cluster.
Clustering.
We first show the justification to represent the training data via their instructions. Specifically, for each dataset, we sample 100 examples and encode their instructions with the all-MiniLM-L6-v2 variant of the Sentence Transformer model [43]. After that, we visualize the data in Figure 3 via t-SNE [51]. According to Figure 3, we observe that (1) samples from different datasets but from the same task are grouped together. For example, coco_vqg, aok_vqg, and ok_vqg are all VQG (visual question generation) tasks (triangle markers) and reside on the left part of the figure. (2) samples from similar tasks are closed to each other. For example, coco_cap and textcaps both belong to the image captioning (CAP) task (small dots) and stay close to each other. Similarly, both the visual question answering (VQA, cross markers) task with datasets (e.g., aok_vqa, coco_vqa, ocrvqa, and ok_vqa) and conversation (CONV) task with datasets (e.g., llava_conversation, llava_detail, and llava_reason) involve answering user questions, which lie in the middle part of Figure 3, suggesting that instructions are good representatives of the training data.
We then cluster all the instructions of the examples in the training data into 64 groups using $k$-means clustering. Figure 4 shows the cluster assignment of the training data. Here, each row in the heatmap denotes a subset of a dataset. The subset is obtained by applying the task template (Sec. 3.1) on the samples of the dataset. We observe (1) different subsets of the same datasets are assigned to similar clusters. For example, aok_vqa, coco_vqa, and ok_vqa are in the first several clusters, while aok_vqg, coco_vqg, and ok_vqg mainly occupy 4 other clusters. (2) datasets of similar tasks are assigned to common clusters. For example, llava_150k including llava_detail, llava_detail and llava_conversation and a series of VQA tasks share the first several clusters as they are to answer questions. Those evidences justify the use of clustering on task instructions as an automatic partition strategy for training datasets.
Routing results.
Figure 5 visualizes the routing decisions of the proposed MoCLE and Sentence-MoLE. We obtain both results from one mixture of LoRA, i.e., one linear module in a layer. The routing results are aggregated by the subset of datasets similar to Figure 4. Shown in Figure 4(a), MoCLE could achieve task-level routing for the inputs. For example, datasets from VQA and VQG tasks are handled by expert 0 and 3 respectively. In addition, VQA expert (i.e., expert 0, which is mostly occupied by VQA task) is fundamental for most of the datasets. That is, most held-out datasets activate the VQA expert. This could be justified by the fact that most vision-language tasks take the form of question answering. However, this does not indicate that the entire model collapses to a single expert because knowledge from other tasks are learned by the universal expert and are leveraged during evaluation. Further, the remaining 3 experts are not dispensable as without them all tasks would clutter in a single expert, resulting in task conflicts.
On the contrary, routing pattern of Sentence-MoLE in Figure 4(b) reveals little correlations between tasks/datasets and experts. That is, different datasets/tasks obtain similar routing decisions, still suffering from task conflicts.
(a) Our MoCLE.
(b) Sentence MoLE.
Figure 5: Routing decisions of one LoRA mixture for MoCLE and Sentence-MoLE. The setup of the vertical axis is similar to Fig. 4 except that we also include the held-out tasks. They are separated by a dotted line on the vertical axis. The horizontal axis corresponds to the index of the LoRA experts.
5 Conclusions
In this paper, we first show that the task conflicts exists in vision language instruction tuning through extensive experiments. To address this, we propose the Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel MoE architecture designed to activate the task-customized model parameters based on the instruction clusters. In addition, we achieve task specialization and generalization in MoCLE simultaneously via a separate universal expert. Comprehensive evaluations of MoCLE on 11 unseen tasks show the effectiveness of the proposed MoCLE method.
Acknowledgments.
We gratefully acknowledge support of MindSpore, CANN (Compute Architecture for Neural Networks) and Ascend AI Processor used for this research.
Appendix
Appendix A Sources of Results in Table 1
For Flamingo-80B [2], Qwen-VL [3], MiniGPT-v2 [5], BLIP-2 [26], and InstructBLIP [11], the results are directly from their papers. For LLaVA and MiniGPT-4, we take the results from [5].
Appendix B Performance on LVLM Benchmarks
We evaluate MoCLE and the reproduced InstructBLIP on several LVLM benchmarks such as POPE [30], MME [13], and MM-Vet [59]. In details, given an image, POPE evaluates the object hallucination of LVLM by determining whether it could correctly respond to the question such as “Is there a broccoli in the image?”. MME and MM-Vet investigate various capacities of LVLM such as perception, cognition, etc. As shown in Table 8, incorporating MoCLE into InstructBLIP significantly reduces the object hallucination. This might result from the fact that questions w.r.t various object categories belong to different instruction clusters [^1] More sample instructions in the clusters are in Sec. H. and are routed to appropriate task experts in MoCLE. In addition, MoCLE also improves various abilities in LVLM as demonstrated in the performance gains in MME and MM-Vet.
Models POPE (adv.) POPE (rand.) POPE (pop.) MME (percep.) MM-Vet
InstructBLIP 77.6 84.3 78.2 1202.9 23.6
+ MoCLE 82.1 88.0 83.7 1222.6 26.0
Table 8: Performance of MoCLE on LVLM benchmarks. We report the F1-score, scores in perceptions, and the overall score for POPE, MME and MM-Vet, where a larger value corresponds to the better performance. POPE (adv.) indicates the adversarial objects splits of POPE. Similarly, POPE (rand.) and POPE (pop.) are for random and popular splits, respectively.
Appendix C Training Details
Following [11], we adopt the same training configurations for the mentioned models such as the proposed MoCLE, the reproduced InstructBLIP, (i.e., InstructBLIP (${}^{{\ddagger}}$) and the task experts in Sec. 1. We train those models with a maximum of 60K steps and a batch size of 128. The AdamW optimizer [23] is used, with $\beta_{1}$ as 0.9, $\beta_{2}$ as 0.999, and a weight decay as 0.05. We apply a linear warmup of the learning rate during the initial 1000 steps, increasing from ${10}^{-8}$ to ${10}^{-5}$, followed by a cosine decay with a minimum learning rate of 0.
Appendix D Weights of the Universal Experts
During training, if some training data obtain a very large weight on a task experts, such data tend to be very specific and might be less beneficial to other tasks. Hence, they get less weight on the universal expert. On the contrary, less specific (a.k.a, more general) data benefit more to other tasks and obtain larger weight on the universal expert. Therefore, the complementarity between the task experts and universal expert achieves good generalization in MoCLE.
Figure 6 shows the average activation weights of the universal experts for different datasets during training. ocr_vqa obtains the lowest weight on the universal expert during training. Indeed, ocr_vqa includes samples that require the model to answer questions such as “What is the title of this book?” and “Who is the author of this book?”. Questions like these have little overlap with the ones in other datasets. However, we observe much higher weights for ok_vqa, ok_vqg, aok_vqa, aok_vqg, llava_*, and coco_vqa. This is consistent with our previous observation in Sec. 4.5 that VQA abilities are fundamental in LVLM.
Figure 6: Weights of the universal expert for different datasets. Colors indicate different datasets.
Appendix E Datasets for Vision-language Instruction Tuning
Here we give a detailed description of the datasets used during instruction tuning. As shown in Table 9, we use 13 datasets in total. Here multiple datasets might be associated with the same data sources because these sources are formatted by different groups of task templates (Sec. F). For LLaVA-Instruct150K [33], we do not apply any task template as it is well formatted. For Web Capfilt, we use COCO+VG+CC+SBU (14M images) as introduced in [25]. In regard to which split is used, please refer to the original InstructBLIP [11] for more details. During training, we sample the datasets using the data balancing strategy proposed in [11].
Datasets Data Source Task Template Group
aok_vqa A-OKVQA [46] VQAMC
aok_vqg A-OKVQA [46] VQG
coco_cap COCO [31] CAP
coco_vqa VQAv2 [15] VQA
coco_vqg VQAv2 [15] VQG
ocrvqa OCR-VQA [40] VQA
ok_vqa OKVQA [39] VQA
ok_vqg OKVQA [39] VQA
textcaps TextCaps [49] OCRCAPS
capfilt Web CapFilt [25] CAP
llava_conversation LLaVA-Instruct150K [33] -
llava_detail LLaVA-Instruct150K [33] -
llava_reason LLaVA-Instruct150K [33] -
Table 9: List of datasets used for training.
Appendix F Task Templates
Following [11], we use the same set of task templates for instruction tuning and zero-shot evaluation. Please refer to Tables 14 and 15 for training and evaluation templates.
Appendix G More Discussion
Limitation.
Although effective, we mainly focus on task conflicts among text-based conversation tasks (e.g., image captioning and VQA) in this paper, while the support of our MoCLE for more complicated visual perception tasks (e.g., object detection [17, 27]) is appealing, which has shown more severe task conflicts with the conversation tasks [68]. Moreover, the applicability of MoCLE to image diffusion models [9, 14, 35, 29] is also a potential research direction, considering the ubiquity of negative transfer phenomenon.
Model responses to counting-related queries. [Uncaptioned image] User How many books are there? New Bing I see that you have sent me an image. According to the description you provided, there are 30 books on the bookshelf. InstructBLIP 100 +MoCLE There are a lot of books on the shelves. It is difficult to say exactly how many books there are, but it appears to be a large number. Table 10: Model responses to counting-related queries.
Model responses to OCR-related queries. [Uncaptioned image] User What are all the exact scene texts in the image? New Bing According to the description you provided, the image contains a license plate with the text “S3302CD”. If you have any other questions or if there is anything else I can help you with, please let me know! InstructBLIP The exact characters in the image are “S 3000 CD.” +MoCLE The exact characters in the image are “S 3302 CD.” Table 11: Model responses to OCR-related queries.
Appendix H Case Studies
In this section, we present several case studies with MoCLE. First, we study its conversation abilities via a range of tasks, including object counting, optical character recognition (OCR), and image introduction. Then we showcase some example instructions sampled from different clusters.
Conversations
In Table 10, we instruct the model to conduct a very difficult object counting task. The correct answer for this question is 63, which is quite hard for existing LVLMs. InstructBLIP failed to give the correct answers, while with MoCLE, InstructBLIP can respond the user query in a much more proper manner.
In Table 11, the model is queried to recognize the character in the image. InstructBLIP performs not so well on this query possibly because OCR-related tasks conflict with other tasks during training. With MoCLE, the model can give correct results.
In Table 12, we ask the model to introduce a famous person in the image. InstructBLIP gives a blunt response to the user query and does not follow the instruction of “introduction”. This might be due to the conflict between image caption and conversation tasks. In the training data, there are a large portion of image caption data that require the model to give a brief description to the image, while the user query in this example expects a detailed introduction to Albert Einstein. With MoCLE, the user query is identified and routed to the correct expert that is specialized at such a conversation task, thus, the model outputs a desired response.
Similarly, in Table 13, we ask the model to describe the image in a detailed manner. InstructBLIP still mistakes this query as an image caption task and gives very short caption to this image. Differently, with MoCLE, the model correctly understands the “in details” in the instruction and provides sufficient details.
Model responses when asked to introduce a famous person. [Uncaptioned image] User Can you give an introduction to this person? New Bing I’m sorry, I cannot give an introduction to this person as I am not allowed to describe people’s identity, gender, race or emotions InstructBLIP albert einstein +MoCLE Albert Einstein was a German-born theoretical physicist who developed the theory of relativity. He is considered one of the most influential scientists of the 20th century. Table 12: Model responses to introduce a celebrity.
Sample Instructions in Clusters
In Table 16, we showcase some sample instructions assigned to different clusters. Though all the instructions in the 4 selected clusters belong to VQA-related tasks, they focus on various perspectives such as food, pet, men, and counting. This justifies the use of a large number of instruction clusters. Moreover, in Table 5, we show that the model performs poorly with only a few clusters, i.e., $K=4$, and gets better as we increase $K$. With a few clusters, VQA, VQG, CONV and CAP tasks would occupy 4 clusters, leaving no more space for samples with different topics. As a result, the model might perform worse when mixing these topics.
Model responses when queried to give detailed image descriptions. [Uncaptioned image] User Introduce me this image in details. New Bing It appears to be a photograph of an adult elephant walking on a beach. The elephant is walking towards the camera and is in the center of the frame. The background consists of the ocean and the sky. The elephant is walking on the sand and there are a few birds in the foreground. The elephant appears to be dark grey in color and has large ears and tusks InstructBLIP an elephant walking along the beach +MoCLE An elephant is walking along the beach, with birds in the background. The elephant’s trunk is extended towards the water, and it appears to be enjoying its time at the beach. Table 13: Model responses to give detailed image description.
Template Group Task Template
CAP A short image caption:
A short image description:
A photo of
An image that shows
Write a short description for the image.
Write a description for the photo.
Provide a description of what is presented in the photo.
Briefly describe the content of the image.
Can you briefly explain what you see in the image?
Could you use a few words to describe what you perceive in the photo?
Please provide a short depiction of the picture.
Using language, provide a short account of the image.
Use a few words to illustrate what is happening in the picture.
VQA {Question}
Question: {Question}
Question A short answer to the question is
Q: {Question} A:
Question: {Question} Short answer:
Given the image, answer the following question with no more than three words. {Question}
Based on the image, respond to this question with a short answer: {Question}. Answer:
Use the provided image to answer the question: {Question} Provide your answer as short as possible:
What is the answer to the following question? “{Question}”
The question “{Question}” can be answered using the image. A short answer is
VQG Given the image, generate a question whose answer is: {Answer}. Question:
Based on the image, provide a question with the answer: {Answer}. Question:
Given the visual representation, create a question for which the answer is “{Answer}”.
From the image provided, craft a question that leads to the reply: {Answer}. Question:
Considering the picture, come up with a question where the answer is: {Answer}.
Taking the image into account, generate an question that has the answer: {Answer}. Question:
[…]
Evaluation Dataset Instruction
GQA, VizWiz, iVQA, MSVD Question: {} Short answer:
NoCaps, Flickr30k A short image description:
TextVQA OCR tokens: {}. Question: {} Short answer:
IconQA Question: {} Options: {}. Short answer:
ScienceQA Context: {} Question: {} Options: {}. Answer:
HatefulMemes This is an image with: “{}” written on it. Is it hateful? Answer:
VSR Based on the image, is this statement true or false? “{}” Answer:
[…]
Table 15: Task templates used during evaluation.
Cluster Instruction Samples Topics
1 “Q: what is being done to the food in the glass fronted box? A:” Food
“Q: what category of pizza would this fall into? Options: (a) vegetarian (b) meat lovers (c) pesto (d) pepperoni, A:”
“what are the large pieces of cake supposed to be?”
“Q: what does this person have on her teeth? Options: (a) braces (b) candy (c) food (d) gum, A:”
“what is the food in? A short answer to the question is”
“what category of pizzas would this be considered?”
2 “Q: what sport is the cartoon dog playing? A:” Pet
“Question: what is likely her favorite animal? Options: (a) cat (b) dog (c) pig (d) sheep, Short answer:”
“Q: what is surrounding the cat? A:”
“Based on the image, respond to this question with a short answer: what color is the cat?. Answer:”
“What might the relationship between the two women and the dog be?”
3 “what type dressing does this man favor?” Men
“Based on the image, respond to this question with a short answer: what are the men doing?. Answer:”
“what is the standing man doing with his arms?”
“what is the man in red shirt doing? Options: (a) laughing (b) crying (c) singing (d) yelling”
“Question: what is the man doing with the pole?”
“Question: why is the man kneeling on the ground?”
4 “how many more animals need to be added to all of these to get the number ten?” Counting
“Question: how many big elephants are inside of this zoo enclosure together? Options: (a) one (b) four (c) two (d) three, Short answer:”
“Q: how many people are seated on the staircase made of wood? A:”
“Question: how many donuts are there?”
“What is the answer to the following question? “how many engines are visible?””
Table 16: Sampled instructions from different clusters.

### v3

@@
 A separate universal expert is further incorporated to improve the generalization capabilities of MoCLE for novel instructions.
-Extensive experiments on 10 zero-shot tasks demonstrate the effectiveness of MoCLE.
+Extensive experiments on 12 zero-shot tasks demonstrate the effectiveness of MoCLE.
 $*{}^{*}$ Equal contribution. ${}^{\dagger}$ Corresponding author.
@@
 During instruction tuning, we only finetune parameters of Q-Former and the LoRA experts while keeping other parts of the LLM frozen.
-Flickr 30K GQA VSR IconQA TextVQA Visdial HM VizWiz SciQA image MSVD QA iVQA Trained using private data Flamingo-80B 67.2 - - - 35.0 - 46.4 31.6 - 35.6 40.7 Qwen-VL (7b) 85.8 $*{}^{*}$ $*{}^{*}$ 35.2 67.1 - - Qwen-VL (7b)-chat 81.0 $*{}^{*}$ $*{}^{*}$ 38.9 68.2 - - Trained using grounding data MiniGPT-v2 (7B) - $*{}^{*}$ 60.6 47.7 - - 58.2 32.9 - - - MiniGPT-v2 (7B)-chat - $*{}^{*}$ 62.9 51.5 - - 58.8 53.6 - - - Trained using public data LLaVA (13B) - 41.3 51.2 43.0 - - - - - - - MiniGPT-4 (13B) - 30.8 41.6 37.6 - - - - - - - BLIP-2 (7B) 74.9 38.6 50.0 39.7 40.1 44.9 50.6 25.3 53.8 18.3 27.5 InstructBLIP (7B) 82.4 49.2 54.3 43.1 50.1 45.2 59.6 34.5 60.5 41.8 52.2 Under the same evaluation InstructBLIP (${}^{{\dagger}}$ 82.9 49.2 65.2 43.1 50.9 44.2 64.3 30.3 59.4 41.5 51.7 InstructBLIP (${}^{{\ddagger}}$ 81.3 48.6 60.8 43.4 53.9 42.7 65.3 29.5 62.0 41.4 51.3 InstructBLIP (${}^{{\ddagger}}$+MoCLE 81.9 (+0.6) 49.3(+0.7) 64.7(+3.9) 46.3(+2.9) 57.1(+3.2) 45.3(+2.6) 65.6(+0.3) 30.4(+0.9) 63.9(+1.9) 42.6(+1.2) 53.2(+1.9)
+Flickr 30K GQA VSR IconQA TextVQA Visdial HM VizWiz SciQA image MSVD QA MSRVTT QA iVQA Trained using private data Flamingo-80B 67.2 - - - 35.0 - 46.4 31.6 - 35.6 17.4 40.7 Qwen-VL (7b) 85.8 $*{}^{*}$ $*{}^{*}$ 35.2 67.1 - - - Qwen-VL (7b)-chat 81.0 $*{}^{*}$ $*{}^{*}$ 38.9 68.2 - - - Trained using grounding data MiniGPT-v2 (7B) - $*{}^{*}$ 60.6 47.7 - - 58.2 32.9 - - - - MiniGPT-v2 (7B)-chat - $*{}^{*}$ 62.9 51.5 - - 58.8 53.6 - - - - Trained using public data LLaVA (13B) - 41.3 51.2 43.0 - - - - - - - - MiniGPT-4 (13B) - 30.8 41.6 37.6 - - - - - - - - BLIP-2 (7B) 74.9 38.6 50.0 39.7 40.1 44.9 50.6 25.3 53.8 18.3 9.2 27.5 InstructBLIP (7B) 82.4 49.2 54.3 43.1 50.1 45.2 59.6 34.5 60.5 41.8 22.1 52.2 Under the same evaluation InstructBLIP (${}^{{\dagger}}$ 82.9 49.2 65.2 43.1 50.9 46.0 64.3 33.7 59.4 41.5 22.7 51.7 InstructBLIP (${}^{{\ddagger}}$ 81.3 48.6 60.8 43.4 53.9 46.3 65.3 34.5 62.0 41.4 23.0 51.3 InstructBLIP (${}^{{\ddagger}}$+MoCLE 81.9 (+0.6) 49.3(+0.7) 64.7(+3.9) 46.3(+2.9) 57.1(+3.2) 46.9(+0.6) 65.6(+0.3) 34.5(+0.0) 63.9(+1.9) 42.6(+1.2) 24.4(+1.4) 53.2(+1.9)
 Table 1: Zero-shot results on the held-out datasets.
@@
 Besides this, we utilize all the other training data, including COCO [31], TextCaps [49], VQAv2 [15], OKVQA [39], A-OKVQA [46], OCR-VQA [40], and LLaVA-Instruct150K [33], for instruction tuning.
-For evaluation datasets, we follow InstructBLIP but only keep Flickr30K [58], GQA [19], VSR [32], IconQA [37], TextVQA [50], Hateful Memes [22], VizWiz [16], ScienceQA [38], MSVD-QA [56], and iVQA [57].
+For evaluation datasets, we follow InstructBLIP but only keep Flickr30K [58], GQA [19], VSR [32], IconQA [37], TextVQA [50], Hateful Memes [22], VizWiz [16], ScienceQA [38], MSVD-QA [56], MSRVTT-QA [56] and iVQA [57].
 InstructBLIP originally uses the NoCaps dataset [1] for the image captioning task, which is not adopted here since we cannot reproduce the reported results, probably due to the incomplete Web CapFilt dataset used.
-In addition, we do not utilize the MSRVTT-QA dataset [56] due to the insufficient computational resources.
 The instructions to conduct all the experiments in this paper are from the original InstructBLIP paper.
@@
 In addition, we achieve task specialization and generalization in MoCLE simultaneously via a separate universal expert.
-Comprehensive evaluations of MoCLE on 11 unseen tasks show the effectiveness of the proposed MoCLE method.
+Comprehensive evaluations of MoCLE on 12 unseen tasks show the effectiveness of the proposed MoCLE method.
 Acknowledgments.
@@
 InstructBLIP failed to give the correct answers, while with MoCLE, InstructBLIP can respond the user query in a much more proper manner.
-In Table 11, the model is queried to recognize the character in the image.
+In Table G, the model is queried to recognize the character in the image.
 InstructBLIP performs not so well on this query possibly because OCR-related tasks conflict with other tasks during training.
@@
 Evaluation Dataset Instruction
-GQA, VizWiz, iVQA, MSVD Question: {} Short answer:
+GQA, VizWiz, iVQA, MSVD, MSRVTT Question: {} Short answer:
 NoCaps, Flickr30k A short image description:

### v4

@@
-Mixture of Cluster-conditional LoRA Experts for Vision-language Instruction Tuning
+Project Page: https://gyhdog99.github.io/projects/mocle/
 Abstract
-Instruction tuning of the Large Vision-language Models (LVLMs) has revolutionized the development of versatile models with zero-shot generalization across a wide range of downstream vision-language tasks.
-However, diversity of training tasks of different sources and formats would lead to inevitable task conflicts, where different tasks conflicts for the same set of model parameters, resulting in sub-optimal instruction-following abilities.
+Instruction tuning of Large Vision-language Models (LVLMs) has revolutionized the development of versatile models with zero-shot generalization across a wide range of downstream vision-language tasks.
+However, the diversity of training tasks of different sources and formats would lead to inevitable task conflicts, where different tasks conflict for the same set of model parameters, resulting in sub-optimal instruction-following abilities.
 To address that, we propose the Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel Mixture of Experts (MoE) architecture designed to activate the task-customized model parameters based on the instruction clusters.
-A separate universal expert is further incorporated to improve the generalization capabilities of MoCLE for novel instructions.
-Extensive experiments on 12 zero-shot tasks demonstrate the effectiveness of MoCLE.
-$*{}^{*}$ Equal contribution. ${}^{\dagger}$ Corresponding author.
+A separate universal expert is further incorporated to improve generalization capabilities of MoCLE for novel instructions.
+Extensive experiments on 11 zero-shot tasks demonstrate the effectiveness of MoCLE.
+Keywords:
+Instruction tuning Mixture of Experts Vision and language
 1 Introduction
-There has been a continuously increasing trend to develop intelligent assistants that can follow human instructions [4, 42, 8], and instruction tuning [55] has become a successful paradigm leveraging large-scale well-formatted instruction data to empower Large Language Models (LLMs) to follow human instructions, demonstrating strong generalization for novel unseen tasks [36].
-Similarly, efforts have been made to introduce similar capabilities to Large Vision-language Models (LVLMs), including LLaVA [33], MiniGPT-4 [67] and InstructBLIP [11], which connects a trainable visual encoder with a frozen LLM and endows the LVLMs with the instruction-following ability by fine-tuning on multi-modal instruction datasets.
-The ability of generalizing to novel unseen instructions necessitates multi-task instruction tuning, i.e., training on a diverse collection of instruction-following tasks, which has been observed in both the LLMs [45, 53, 10] and LVLMs [3, 63, 28].
-However, complexity of various instruction tasks brings difficulties for the model fine-tuning.
-Specifically, Wei et al.
-[55] observe that under certain model sizes, multi-task instruction tuning even fails to bring performance gains for zero-shot tasks compared to the original models, mainly attributed to the negative transfer phenomenon [62] during multi-task instruction tuning, where the model struggles to optimize the losses of multiple conflicted tasks, leading to sub-optimal performance for these tasks.
+There has been a continuously increasing trend to develop intelligent assistants that can follow human instructions [4, 48, 8], with instruction tuning emerging as a notably effective approach.
+This method leverages large-scale well-formatted instruction data to empower Large Language Models (LLMs) to execute various human instructions, showcasing their ability to generalize across novel unseen tasks [41].
+Likewise, efforts have been made to introduce similar capabilities to Large Vision-language Models (LVLMs) [3, 66, 63, 10, 5], including LLaVA series [38, 37], MiniGPT-4 [73] and InstructBLIP [14].
+They connect a trainable visual encoder with a frozen LLM and endow the LVLMs with the instruction-following ability by finetuning on diverse multi-modal instruction datasets.
+It is observed in both the LLMs [51, 59, 13] and LVLMs [3, 69, 31] that the ability to generalize to novel unseen instructions necessitates multi-task instruction tuning, i.e., training on a diverse collection of instruction-following tasks.
+However, the complexity of various instruction tasks brings difficulties for the model fine-tuning.
+Specifically, [60] find that under certain model sizes, multi-task instruction tuning even fails to bring performance gains for zero-shot tasks compared to the original models, mainly attributed to the negative transfer phenomenon [68] during multi-task instruction tuning, where the model struggles to optimize the losses of multiple conflicted tasks, leading towards the sub-optimal experimental performance for these tasks.
 Figure 1: Performance of the instruction-finetuned LVLMs on zero-shot tasks, where larger values indicate better performance.
-Only 3 out of 8 tasks benefit from instruction tuning from all the data, while the task experts show better performance on the other 5 tasks (i.e., Flickr 30K, HM, ScienceQA, IconQA, and VizWiz).
+Only 2 out of 7 tasks benefit from instruction tuning from all the data, while the task experts show better performance on the other 5 tasks (i.e., Flickr 30K, GQA, HM, SciQA and IconQA).
 Similarly, tasks for vision-language instruction tuning (e.g., visual question answering and image captioning) focus on different perspectives of LVLMs, and thus, resulting in conflicts when sharing all the parameters as most studies adopt.
-In our preliminary studies shown in Figure 1, we first split the instruction data into two disjoint subsets, including “cap” for image captioning and “vqa” for visual question answering.
+In our preliminary study, we first split the instruction data into two disjoint subsets, including “cap” for image captioning and “vqa” for visual question answering.
 We then train InstructBLIP [14] using LoRA [21] on 3 kinds of data: “cap”, “vqa” and “full” (the full data) separately, to obtain 3 set of parameters (i.e., task experts).
-Following the held-out evaluation protocol in [11], we evaluate these experts on the unseen datasets/tasks with the best expert.
-As shown in Figure 1, on 5 out of the 8 downstream tasks, InstructBLIP instruction-tuned on all the data is outperformed by the task expert that is finetuned on only a subset of data.
-Among the 5 tasks, Flickr30k belongs to “cap”, and SciQA, IconQA and VizWiz belong to “vqa”, revealing that instruction tuning on the similar tasks brings positive transfer to downstream tasks, while training on the full data with dissimilar tasks could hurt generalization performance.
-The task experts above is a naive solution to negative transfer, where we manually partition the training tasks and train each expert individually.
-However, there are several limitations of this strategy, (1) The taxonomies such as “vqa” and “cap” require human expertise and become difficult to scale as the number of tasks grows; (2) The ability to generalize to unseen tasks is inhibited, as we do not know which expert to choose for novel tasks and some new tasks might benefit from multiple training tasks (e.g., VSR and TextVQA as in Figure 1).
-In this regard, specialization and generalization of LMMs becomes a dilemma.
-Our work aims to develop an automatic and practical partition strategy as well as a network architecture that can strike a balance between specialization and generalization.
+Following the held-out evaluation protocol [14], we evaluate these experts on the unseen datasets/tasks with the best expert.
+As shown in Figure 1, on 5 out of the 7 downstream tasks, the InstructBLIP instruction-tuned on all the data is outperformed by the task expert finetuned with only a subset of data.
+Among the 5 tasks, the Flickr30k belongs to “cap”, and SciQA, GQA and IconQA belong to “vqa”, revealing that instruction tuning on the similar tasks brings positive transfer to downstream tasks, while training on the full data with dissimilar tasks could hurt the generalization performance.
+The disjoint task expert above is a naïve solution to negative transfer, where we manually partition training tasks and train each expert individually.
+However, there are several limitations of this strategy: (1) The taxonomies such as “vqa” and “cap” require human expertise and become difficult to scale as the number of tasks grows.
+(2) The ability to generalize to unseen tasks is inhibited, as we do not know which expert to choose for novel tasks, while some new tasks might benefit from multiple training tasks (e.g., VSR and TextVQA as in Figure 1).
+In this regard, specialization and generalization of LVLMs becomes a dilemma.
+Our work aims to develop an automatic and practical partition strategy and a network architecture that can strike a balance between specialization and generalization.
 We propose Mixture of Cluster-conditional LoRA Experts (MoCLE) for vision-language instruction tuning.
-Under the proposed MoCLE framework, we first cluster instructions of all the training data into several centroids via a pretrained clustering model.
-In this way, similar tasks that could bring positive transfer to each other are grouped together in a cluster (more justification for the use of instruction clusters is detailed in Sec.
+Under the proposed MoCLE framework, we first cluster instructions of all the training data into several centroids via a pre-trained clustering model.
+In this way, similar tasks that could bring positive transfer to each other are automatically grouped into the same cluster, and different tasks that may cause conflict are separated (more justification for the usage of instruction clusters is detailed in Sec.
 3.3).
-Then using the cluster as a condition, at each layer of the model, a router dispatches the input data to one of the specialized task experts and also a universal expert that is shared among all the data.
-Finally, these two set of outputs are combined and passed to the next layer.
+Then we construct several task experts with each concentrating on the specific cluster.
+Using the cluster as a condition, a router dispatches the input data to one of the specialized task experts and a universal expert that is shared among all the data.
 As we activate a specialized expert for a group of similar tasks, tasks that are less alike are learned via separate experts, mitigating task conflicts.
 Meanwhile, since the universal expert trained on all tasks also contributes to the model outputs, we can enjoy generalization and specialization simultaneously.
-We validate the effectiveness of MoCLE on the state-of-the-art InstructBLIP [11].
-MoCLE achieves remarkable improvement on the held-out tasks [11] including image captioning (e.g., Flickr30K [58]), text-based visual question answering (e.g., TextVQA [50]), knowledge-based visual question answering (e.g., ScienceQA [38]) and novel visual reasoning (e.g., Visual Spatial Reasoning [32]).
+We validate effectiveness of MoCLE on the state-of-the-art InstructBLIP [14].
+MoCLE achieves significant improvement on the held-out tasks [14] including the image captioning (e.g., Flickr30K [64]), the text-based visual question answering (e.g., TextVQA [56]), the knowledge-based visual question answering (e.g., ScienceQA [42]), and novel visual reasoning (e.g., Visual Spatial Reasoning [36]).
 The main contributions of this work contain three parts.
   1. 1.
-    We identify the negative transfer phenomenon [34, 65] of task conflicts during instruction tuning of LVLMs.
+    We identify the negative transfer phenomenon[40, 71] of task conflicts during the instruction tuning of LVLMs.
   2. 2.
-    We propose the Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel parameter-efficient finetuning framework suitable for the vision-language instruction tuning, to mitigate task conflicts and enjoy the benefits of huge data training simultaneously.
+    We propose Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel parameter-efficient finetuning framework suitable for the vision-language instruction tuning, to mitigate the task conflicts and enjoy the benefits of huge data training simultaneously.
   3. 3.
-    The proposed MoCLE method achieves remarkable performance gains on the unseen held-out tasks compared with a strong baseline model InstructBLIP [11].
+    Our proposed MoCLE achieves remarkable performance gains on the unseen held-out tasks compared with a strong baseline model InstructBLIP [14].
 2 Related Work
-Multi-task Instruction Tuning.
+2.0.1 Multi-task Instruction Tuning.
 Instruction tuning, as elucidated in [51, 60], is a technique where a language model is finetuned over a collection of tasks and instruction templates that tells the model the intention of the task.
-The primary goal of this technique is to teach the model to understand the relationship between the instruction and the input/output pairs, so that the model could generalize to unseen task given new instructions.
+The primary goal of this technique is to teach the model to understand the relationship between the instruction and the input/output pairs, so that the model could generalize to unseen task given novel instructions.
 Increasing the number of diverse instructions [51], the number of tasks [59, 13], and diversity of data [72] have been shown to improve the performance.
-However, Wei et al.
-[55] observed under certain model sizes, instruction tuning of language models fail to bring performance gains to unseen tasks when compared with untuned base models.
+However, [60] observe under certain model sizes, instruction tuning of language models fail to bring performance gains to unseen tasks when compared with untuned base models.
 They argue this is due to the fact all model capacity is used to learn the mixture of instruction tuning tasks.
-In this work, we aim to mitigate such issue in vision-language instruction tuning by specialized experts.
-Large Multi-modal Models.
+In this work, we aim to mitigate such issue in vision-language instruction tuning by the specialized experts.
+2.0.2 Large Multi-modal Models.
 To inherit strong reasoning and generalization abilities of LLMs and meanwhile understand visual inputs, recent works including Flamingo [2], BLIP-2 [28], LLaMA-adapter [67] connect LLMs to the visual modality with a frozen visual encoder and a trainable module, which is then finetuned with plain image-text pairs.
-To further improve instruction following abilities and response quality, LLaVA finetunes the whole LLM on vision-language conversational data generated by GPT-4.
-MiniGPT-4 performs training on ChatGPT-generated image captions instead of traditional caption data that are much shorter.
-InstructBLIP collects a wide range of academic VQA datasets and manually crafts 10 to 15 distinct instruction templates for each kind of task.
-In this paper, the proposed MoCLE is built on InstructBLIP as the base model.
+To further improve instruction following abilities and response quality, LLaVA series [38, 37] finetune the whole LLM on the vision-language conversational data generated by GPT-4.
+MiniGPT-4 [73] performs training on ChatGPT-generated image captions instead of traditional caption data that are shorter and informational sparser.
+InstructBLIP [14] collects a wide range of academic VQA datasets and manually crafts 10 to 15 distinct instruction templates for each kind of task.
+In this paper, we primarily consider InstructBLIP as the base model.
+2.0.3 Mixture of Experts.
+MoE models [23, 24, 53] are considered as an effective way to increase the model capacity in terms of parameter size.
+Within MoEs, certain parts of the model are activated while the computation is kept the same or close to its dense counterpart.
+Recently, it has been thoroughly investigated in the field of computer vision[50, 71], natural language processing[27, 15] and multi-modal learning[47, 54, 12, 11, 34].
+MoCE [71] tackles negative transfer in the self-supervised visual pre-training[6, 7] using a mixture of cluster experts.
+Different from MoCE requiring to be finetuned on downstream tasks, our MoCLE studies the zero-shot generalization to unseen tasks.
+Another line of works combines the MoE with adapters.
+AdaMix[58] randomly selects adapters during training and adopts the average of expert parameters during inference.
+Zadouri et al.
+[65] propose mixture of LoRAs for instruction tuning of LLMs.
+Concurrently, MoE-LLaVA [34] sparsifies LVLMs via MoEs for more efficient model scaling.
+The most relevant to ours include the Octavius [12] and LLaVA-MoLE [11], which also address task conflicts in vision-language instruction tuning by applying the MoE architecture with LoRAs.
+However, they do not explore the issue of instruction generalization in the architecture designs.
+Instead, we design a novel Mixture of Cluster-conditional LoRA Experts together with the universal experts, which is the first work addressing both task conflicts and generalization with a unified architecture.
+3 Methodology
+In this section, we start with the formulation of LVLM instruction tuning and the analysis of the limitations of task experts.
+We then introduce the proposed MoCLE in detail, with the overall framework shown in Figure 2.
 Figure 2: Overall pipeline of MoCLE.
-The left part is to cluster training data (i.e., a collection of vision-language tasks with diverse instructions) into several groups based on their instructions via $k$-means clustering algorithm.
+The left part is to cluster training data (a collection of vision-language tasks with diverse instructions) into several groups based on the instructions via $k$-means clustering algorithm.
 On the right is the MoCLE framework.
-$\mathbf{x}$ are the hidden representations of the sample (e.g., a sequence of tokens for images, instructions and the raw text inputs in the LLM).
-The input $\mathbf{x}$ is handled by 3 modules, including the universal expert, the task expert with the largest gating value, and the linear module of the LLM.
+$\mathbf{x}$ is the hidden representation of the sample (e.g., a sequence of tokens for images, instructions and the raw text inputs in the LLM) which is handled by 3 modules, including the universal expert, task expert, and the original linear module of the LLM.
 After that, the outputs of these modules are aggregated and passed to the next layer.
 Only experts, the cluster embedding and the gate parameters are updated during training, while parameters of the LLM are frozen.
-We insert such structure into the query and value projection of self-attention blocks of each layer by default.
-Mixture of Experts.
-MoE models [20, 21, 47] are considered as an effective way to increasing the model capacity in terms of parameter size.
-In MoE, certain parts of the model are activated while the computation is kept the same or close to its dense counterpart.
-Recently, it has been thoroughly investigated in the field of computer vision [44, 65], natural language processing [24, 12] and multi-modal learning [41, 48].
-Among those studies, MoCE [65] tackles negative transfer in the self-supervised visual pre-training [6, 7] using a mixture of cluster experts.
-Different from MoCE that needs to be finetuned on downstream tasks, the proposed MoCLE method studies the zero-shot generalization to unseen tasks.
-Another line of works combines MoE with adapters.
-AdaMix [54] randomly selects adapters during training and uses the average parameters of those experts during the inference.
-Zadouri et al.
-[60] proposes mixture of LoRAs for instruction tuning of LLMs, achieving comparable results with the fully-finetuned counterpart.
-However, neither of them explicitly addresses task conflicts and generalization to unseen tasks, which is the focus of our work.
-3 Methodology
-In this section, we start with the formulation of LVLM instruction tuning and the analysis of the limitations of task experts.
-We then introduce the proposed MoCLE in detail.
-Figure 2 shows the overall framework of MoCLE.
+We insert such structure into the query and value projection in self-attention blocks of each layer by default.
 3.1 Problem Formulation
-Suppose that there are a set of datasets which are divided into held-in and held-out datasets, a large vision-language model is first fine-tuned on the held-in and then evaluated in a zero-shot manner on the held-out datasets.
-To unify input-output formats and promotes generalization in instruction tuning, several task templates $\{T_{i}\}$ are designed to wrap the raw inputs, i.e., a pair of text and image $(X_{\text{txt}},X_{\text{img}})$ from the corresponding dataset, where $T_{i}$ denotes the $i$-th template.
+Suppose that there is a set of datasets that are divided into held-in and held-out datasets [14], where a large vision-language model is first finetuned on the held-in and then evaluated on the held-out datasets in a zero-shot manner.
+To unify and diversify input-output formats and promotes in instruction tuning, several task templates $\{T_{i}\}$ are designed to wrap the raw inputs, a pair of text and image $(X_{\text{txt}},X_{\text{img}})$ from the corresponding dataset, where $T_{i}$ denotes the $i$-th template.
 For example, “Given the image, answer the question with no more than three words.
-{Question}” is a template for the visual question answering tasks.
-An instruction $I\equiv T_{i}(X_{\text{txt}})$ wraps the text input using the template.
+{Question}” is a template for visual question answering tasks.
+The instruction is defined as $I\equiv T_{i}(X_{\text{txt}})$ that wraps text inputs using the template.
 3.2 Limitation of Task Experts
-As observed in Figure 1, allocating task experts based on the categories of the training datasets brings positive transfer to similar downstream tasks.
-However, there are several limitations of this solution.
-The first limitation is that taxonomies such as “vqa” and “cap” require delicate human designs with expertise in the task performed.
-The second limitation is that when a new task comes, we do not know which expert to use.
+Figure 1 indicates that allocating task experts based on the categories of training datasets brings positive transfer to similar downstream tasks.
+Nevertheless, this approach has several notable limitations.
+First, taxonomies such as “vqa” and “cap” require delicate human designs with expertise in the task performed.
+Another limitation is the challenge of selecting the appropriate expert for a newly introduced task when testing.
 Further, with tasks trained in an isolated manner, knowledge sharing among different tasks is prevented, compromising generalization to unseen tasks and instructions as some might benefit from multiple tasks.
-Hence, we propose the MoCLE in the following sections to alleviate these two limitations.
+Hence, we propose MoCLE in the following sections to address these limitations.
 3.3 Clustering Data by Instructions
-The goals of partitioning the training data are two-fold.
-We target to train a task expert within a collection of similar tasks so as to prevent task conflicts.
-Meanwhile, we expect new tasks to be assigned to the proper experts based on its cluster without manual intervention.
+The purposes of partitioning the training data are two-fold.
+We target at training a task expert within a collection of similar tasks so as to prevent task conflicts.
+Meanwhile, we expect novel tasks to be assigned to the proper experts based on their cluster without manual intervention.
 To achieve these, we should find a proper description for the task.
@@
 The raw inputs further infuse this framework with the substantive content that the model must process and respond to.
-Therefore, by conducting clustering on the instructions, we can partition training data without human labor into several clusters, each containing a set of similar tasks.
-We further verify this through experiments later in Sec.
+Therefore, by conducting clustering on the instructions, we can partition training data without human labor into several clusters, each containing a set of similar tasks, which is further verified through experiments in Sec.
 4.5.
-Moreover, even when novel instruction tasks come, we can obtain a cluster index for them using the pretrained clustering model.
-Formally, denoting $\mathcal{E}(\cdot)$ as a pretrained sentence encoder, $\mathbf{e}_{i}=\mathcal{E}(I_{i})$ is the sentence representation of an instruction $I_{i}$, we use the $k$-means clustering algorithm to group all the instructions in the training datasets into $K$ clusters by iteratively minimizing $\sum_{j=1}^{K}\sum_{\mathbf{e}_{i}\in S_{j}}\|\mathbf{e}_{i}-\mathbf{c}_{j}\|^{2}$, where $S_{j}$ denotes the set of instructions assigned to the $j$-th cluster and $\mathbf{c}_{j}$ denotes the centroid of the $j$-th cluster.
-In each iteration of the $k$-means clustering algorithm, each instruction is assigned to the nearest centroid with all centroids updated to the average of instructions in the corresponding cluster.
-3.4 Mixture of Cluster-conditional Experts
-In Sec.
-1, we defined a static one-to-one mapping between the task categories and expert parameters.
-For example, the data in the $i$-th cluster are processed by expert-$j$ for any layer at all the time.
-Such static mapping is sub-optimal since $N$ tasks categories require $N$ task experts, resulting prohibited model sizes as $N$ becomes large.
-In fact, an expert does not necessarily have to handle only one kind of tasks all the time.
-That is, at different layers, a task could have different choices of experts.
-Concretely, consider $L$ layers each with $N$ experts, then there are $N^{L}$ number of choices in total.
-This would accommodate much more tasks while keep $N$ small.
-Yet, under such a circumstance, how to choose the expert for a given task remains unclear.
+Moreover, even when novel instruction tasks come, we can obtain a cluster index for them using the pre-trained clustering model without human intervention.
+Formally, denoting $\mathcal{E}(\cdot)$ as a pre-trained sentence encoder, $\mathbf{e}_{i}=\mathcal{E}(I_{i})$ is the sentence representation of an instruction $I_{i}$, we use the $k$-means clustering algorithm to group all instructions in the training datasets into $K$ clusters by iteratively minimizing $\sum_{j=1}^{K}\sum_{\mathbf{e}_{i}\in S_{j}}\|\mathbf{e}_{i}-\mathbf{c}_{j}\|^{2}$, where $S_{j}$ denotes the set of instructions assigned to the $j$-th cluster and $\mathbf{c}_{j}$ denotes the centroid of the $j$-th cluster.
+In each iteration of the $k$-means clustering algorithm, each instruction is assigned to the nearest centroid with all centroids updated as the average of the instruction representations in the corresponding cluster.
+3.4 Mixture of Cluster-conditional LoRA Experts
+In addition to considerations at the data level, we also suggest an architectural design to tackle the issue of negative transfer.
 We propose Mixture of Cluster-conditional LoRA Experts (MoCLE) that learns to activate the LoRA expert at each layer given the cluster of the data.
-Specifically, denoting by $E$ the number of experts, we introduce a gate vector $\mathbf{G}\in\mathbb{R}^{E}$ that, given any input $\mathbf{x}_{i}$, determines the experts to which the input is routed.
+Specifically, denoting $E$ as the number of experts, we introduce a gate vector $\mathbf{G}\in\mathbb{R}^{E}$ that, given any input $\mathbf{x}_{i}$, determines the experts to which the input is routed.
 The gate vector is calculated as,
@@
 This is the key for the model to choose proper task experts for the input data.
-To endow the clustering embedding with task information, we initialize it to be the centroid of the represented cluster.
+To endow the clustering embedding with task information, we initialize it to be the centroid of the corresponding cluster.
 $\mathbf{W}_{\text{gate}}$ denotes trainable weights of the linear gate, $\boldsymbol{\epsilon}\sim N(0,\frac{1}{E})$ denotes a noise term that is designed to add some randomness to the expert choosing process, which encourages MoCLE to explore multiple combinations of experts during training.
 $\tau$ denotes a temperature hyperparamter.
-The output $\mathbf{y}_{i}$ is then computed as a weighted sum of the output of each expert on the input $\mathbf{x}_{i}$ as,
-$\mathbf{y}_{i}=\sum_{e=1}^{E}G_{e}\mathbf{W}_{e}\mathbf{x}_{i},$ (2)
-where $\mathbf{W}_{e}$ represents the linear projection weights of the $e$-th expert, and $G_{e}$, the $e$-th entry in $\mathbf{G}$, indicates the contribution of the $e$-th expert to $\mathbf{y}_{i}$.
+The output $\mathbf{y}_{i}$ is then computed as a sum of the weighted output of each expert and the original LLM linear layer [21] on the input $\mathbf{x}_{i}$ as,
+$\mathbf{y}_{i}=\sum_{e=1}^{E}G_{e}\mathbf{W}_{e}\mathbf{x}_{i}+\mathbf{W}_{0}\mathbf{x}_{i},$ (2)
+where $\mathbf{W}_{0}$ denotes the pre-trained linear layer of LVLMs, $\mathbf{W}_{e}$ is the linear projection weights of the $e$-th LoRA expert, and $G_{e}$, the $e$-th entry in $\mathbf{G}$, indicates the contribution of the $e$-th expert.
 In this formulation, the model learns to choose the expert that is suitable for the data by examining its cluster.
-In addition, as a task has different choices of expert at different layers, expert number scales well with the number of task clusters.
-We mitigate task conflicts with only a few experts.
+We will show in Sec.
+4.4.3 that the condition of clusters in expert routing is indispensable for task generalization with significant superiority over typical routing conditions including sentence embeddings and token representations.
 3.5 Universal Expert
@@
 3.4 still hurts the generalization ability of the entire model.
-We hypothesize that this might be attributed to the fact that instruction-tuned models generalize to unseen tasks by training on many instructions, while in our case, each expert sees fewer instructions than the original dense model does.
+We hypothesize that this might be attributed to the fact that the instruction-tuned models generalize to unseen tasks by training on extensive instructions [60], while in our formulation, each expert sees fewer instructions than the original dense model.
 Hence, we propose a universal expert that could learn from all the training data.
@@
 Then the output for all the experts is expressed as,
-$\mathbf{y}_{i}=\left(\sum_{e=1}^{E}G_{e}\mathbf{W}_{e}+(1-G_{\text{max}})\mathbf{W}_{u}\right)\mathbf{x}_{i}.$ (3)
+$\mathbf{y}_{i}=\left(\sum_{e=1}^{E}G_{e}\mathbf{W}_{e}+(1-G_{\text{max}})\mathbf{W}_{u}\right)\mathbf{x}_{i}+\mathbf{W}_{0}\mathbf{x}_{i}.$ (3)
 In this formulation, we additionally train a universal expert parameterized by $\mathbf{W}_{u}$.
 Different from task experts that are activated only for specific model inputs, the universal expert is activated for any inputs.
-The final output is the weighted sum of outputs of one of the experts and the universal expert.
+The final output is the weighted sum of outputs from one of the experts and the universal expert puls the original LVLM’s output.
 The weighting for this universal expert is the complement of that of the chosen task expert.
 Consequently, the task expert learns distinct skills for certain tasks while the universal expert masters holistic understanding of the training corpus.
-The synergy between them offers both specialization and generalization for the LVLM.
+The synergy between them offers both specialization and generalization for the LVLMs with MoCLE.
 4 Experiment
-In this section, we empirically evaluate MoCLE on various downstream tasks in a zero-shot manner.
+In this section, we conduct an assessment of MoCLE across multiple downstream tasks under a zero-shot setting.
+Initially, we detail the experimental settings and implementation details, which is followed by a description of the datasets and instructions employed, along with the outcomes of our evaluations.
+Lastly, we present an ablation study and visualizations of clustering and routing results.
 4.1 Settings and Implementation Details
-In all the experiments, we utilize the InstructBLIP [11] model, particularly its Vicunna-7B variant [64], as the large vision-language model (LVLM) to assess the effectiveness of the proposed method.
-We encode the instructions of different datasets using the all-MiniLM-L6-v2 variant of the Sentence Transformer model [43] and cluster their embeddings into $64$ clusters via $k$-means clustering algorithm.
-During the vision-language instruction tuning, we initialize the model based on pre-trained InstructBLIP checkpoints [11] and apply the proposed MoCLE on the LLM (i.e., Vicunna-7B).
+In all the experiments, we utilize the InstructBLIP[14] model, particularly its Vicuna-7B variant [70], as the base LVLM to assess the effectiveness of MoCLE.
+We encode all the instructions of different datasets using the all-MiniLM-L6-v2 variant of the Sentence Transformer model [49] and cluster their embeddings into $64$ clusters via $k$-means clustering algorithm.
+We apply the proposed MoCLE on the LLM (i.e., Vicuna-7B) of InstructBLIP.
 Specifically, for each transformer layer in the LLM, we insert the MoCLE, which consists of 4 task LoRA experts and one LoRA universal expert, into the query and value projection of self-attention blocks.
@@
 During instruction tuning, we only finetune parameters of Q-Former and the LoRA experts while keeping other parts of the LLM frozen.
-Flickr 30K GQA VSR IconQA TextVQA Visdial HM VizWiz SciQA image MSVD QA MSRVTT QA iVQA Trained using private data Flamingo-80B 67.2 - - - 35.0 - 46.4 31.6 - 35.6 17.4 40.7 Qwen-VL (7b) 85.8 $*{}^{*}$ $*{}^{*}$ 35.2 67.1 - - - Qwen-VL (7b)-chat 81.0 $*{}^{*}$ $*{}^{*}$ 38.9 68.2 - - - Trained using grounding data MiniGPT-v2 (7B) - $*{}^{*}$ 60.6 47.7 - - 58.2 32.9 - - - - MiniGPT-v2 (7B)-chat - $*{}^{*}$ 62.9 51.5 - - 58.8 53.6 - - - - Trained using public data LLaVA (13B) - 41.3 51.2 43.0 - - - - - - - - MiniGPT-4 (13B) - 30.8 41.6 37.6 - - - - - - - - BLIP-2 (7B) 74.9 38.6 50.0 39.7 40.1 44.9 50.6 25.3 53.8 18.3 9.2 27.5 InstructBLIP (7B) 82.4 49.2 54.3 43.1 50.1 45.2 59.6 34.5 60.5 41.8 22.1 52.2 Under the same evaluation InstructBLIP (${}^{{\dagger}}$ 82.9 49.2 65.2 43.1 50.9 46.0 64.3 33.7 59.4 41.5 22.7 51.7 InstructBLIP (${}^{{\ddagger}}$ 81.3 48.6 60.8 43.4 53.9 46.3 65.3 34.5 62.0 41.4 23.0 51.3 InstructBLIP (${}^{{\ddagger}}$+MoCLE 81.9 (+0.6) 49.3(+0.7) 64.7(+3.9) 46.3(+2.9) 57.1(+3.2) 46.9(+0.6) 65.6(+0.3) 34.5(+0.0) 63.9(+1.9) 42.6(+1.2) 24.4(+1.4) 53.2(+1.9)
-Table 1: Zero-shot results on the held-out datasets.
-Here, Visdial, HM and SciQA denote Visual Dialog, HatefulMemes and ScienceQA, respectively.
-For ScienceQA, we only evaluate the set with image context.
-Following previous works [11], we report the CIDEr score [52] for Flickr30K, accuracy for iVQA, AUC score for HatefulMemes, and Mean Reciprocal Rank (MRR) for Visual Dialog.
-For all other datasets, we report top-1 accuracy (%).
-Notice results with * means the training split of the evaluation dataset is used during instruction tuning, thus not zero-shot.
-Best results obtained by models trained with public data are marked in bold and the second best are underlined.
-Superscript ${}^{{\dagger}}$ denotes the official InstructBLIP checkpoint.
-Superscript ${}^{{\ddagger}}$ denotes our reproduced version, into which MoCLE is incorporated.
+Models GQA VSR IconQA Visdial MME POPE A-OKVQA OKVQA VQAv2 Direct MC (test) (test-dev) LLaVA (13B) 41.3 51.2 43.0 - - 67.0 50.5 - 54.4 - MiniGPT-4 (13B) 30.8 41.6 37.6 - 581.7 71.4 38.7 - 37.5 - InstructBLIP (7B) 48.6 60.8 43.4 46.3 1202.9 77.6 58.8 73.8 57.0 77.4 49.3 64.7 46.3 46.9 1222.6 82.1 61.5 78.2 59.8 78.9 MoCLE (7B) (+0.7) (+3.9) (+2.9) (+0.6) (+19.7) (+4.5) (+2.7) (+4.4) (+2.8) (+1.5)
+Table 1: Zero-shot results on the held-out datasets (GQA, VSR, IconQA, Visdial) and LVLM benchmarks (MME, POPE) and evaluation on held-in datasets (A-OKVQA, OKVQA, VQAv2).
+Here Direct and MC denote directly answering and multiple choices.
+Best results are marked in bold.
+Models Flickr 30K TextVQA HM SciQA image MSVD QA MSRVTT QA iVQA InstructBLIP (7B) 81.3 53.9 65.3 62.0 41.4 23.0 51.3 81.9 57.1 65.6 63.9 42.6 24.4 53.2 MoCLE (7B) (+0.6) (+3.2) (+0.3) (+1.9) (+1.2) (+1.4) (+1.9)
+Table 2: Zero-shot results on the held-out datasets. Here, HM and SciQA denote HatefulMemes and ScienceQA, respectively.
 4.2 Datasets and Instructions
@@
 Besides this, we utilize all the other training data, including COCO [35], TextCaps [55], VQAv2 [19], OKVQA [45], A-OKVQA [52], OCR-VQA [46], and LLaVA-Instruct150K [38], for instruction tuning.
-For evaluation datasets, we follow InstructBLIP but only keep Flickr30K [58], GQA [19], VSR [32], IconQA [37], TextVQA [50], Hateful Memes [22], VizWiz [16], ScienceQA [38], MSVD-QA [56], MSRVTT-QA [56] and iVQA [57].
-InstructBLIP originally uses the NoCaps dataset [1] for the image captioning task, which is not adopted here since we cannot reproduce the reported results, probably due to the incomplete Web CapFilt dataset used.
+For the zero-shot evaluation datasets, we follow InstructBLIP but only keep Flickr30K [64], GQA [22], VSR [36], IconQA [43], TextVQA [56], Hateful Memes [25], ScienceQA [42], MSVD-QA [61], MSRVTT-QA [61] and iVQA [62].
+NoCaps dataset [1] is originally utilized for the image captioning task, which, however, is not adopted here since we cannot reproduce the reported results, probably due to the incomplete Web CapFilt dataset used.
+In addition, we consider another two widely adopted LVLM benchmarks, i.e., MME [16] and POPE [33] for zero-shot evaluations.
+For held-in evaluation, we consider the VQAv2 [19], OKVQA [45] and A-OKVQA [52].
+For ScienceQA, we only evaluate the set with image context.
+Following previous works [14], we report the CIDEr score [57] for Flickr30K, the iVQA accuracy for iVQA, AUC score for HatefulMemes, and Mean Reciprocal Rank (MRR) for Visual Dialog.
+For LVLM benchmarks, we report the perception score for MME and F1 score for the adversarial split of POPE.
+For all other datasets, we report the top-1 accuracy (%).
 The instructions to conduct all the experiments in this paper are from the original InstructBLIP paper.
 The detailed introductions of the held-in and held-out tasks/datasets are also presented in [14].
-4.3 Zero-shot Evaluations
-Table 1 presents the experimental results on multiple zero-shot vision-language tasks.
-We build our method upon the Vicunna-7B variant of InstructBLIP.
-The InstructBLIP (${}^{{\ddagger}}$ is our reproduced version of InstructBLIP.
+4.3 Evaluation Results
+Table 1 and Table 2 present the experimental results on multiple zero-shot and held-in vision-language tasks.
+We build our method upon the Vicuna-7B variant of InstructBLIP.
 Compared to this model, the proposed MoCLE method shows considerable performance improvement.
-Specifically, on conventional tasks such as visual question answering (VQA), we observe a 2.9% improvement on IconQA.
-On tasks with novel formats such as Visual Spatial Reasoning (VSR), TextVQA and ScienceQA image, we obtain an absolute performance gain of 3.9%, 3.2%, and 1.9%, respectively.
-This indicates that the proposed MoCLE method facilitates generalization to unseen tasks/instructions since the instructions of VSR, TextVQA and ScienceQA image are unseen during training.
-In addition, the proposed MoCLE method performs well when compared with existing popular LVLMs including LLaVA (13B) and MiniGPT-4 (13B).
-For example, on IconQA, the proposed MoCLE method outperforms LLaVA (13B) and MiniGPT-4 (13B) by 3.3% and 8.7%, respectively.
-In summary, the proposed MoCLE model obtains best results on IconQA, TextVQA, HM, and ScienceQA image across models trained using public data.
+Specifically, on held-out datasets such as IconQA, Visual Spatial Reasoning (VSR), TextVQA and ScienceQA datasets, we obtain an absolute performance gain of 2.9% 3.9%, 3.2%, and 1.9%, respectively.
+This indicates that the proposed MoCLE method facilitates generalization to unseen tasks/instructions since these datasets are unseen during training.
+In addition, we observe considerable improvement on common LVLM benchmarks (i.e., MME and POPE).
+POPE evaluates whether the model suffers from object hallucinations while MME considers a wide range of abilities such as math, OCR, color and understanding of arts.
+Therefore, our MoCLE cannot only promote task generation, but also improve overall abilities of the LVLM.
+Further, the held-in tasks whose datasets are seen during training could also benefit from the proposed MoCLE.
+Specifically, with our MoCLE, an absolute improvement of 4.4%, 2.8% and 1.5% could be observed on A-OKVQA (MC), OKVQA and VQAv2.
+This indicates that MoCLE could effectively alleviate the task conflicts during multi-task learning.
 4.4 Ablation Studies
-In this section, we first ablate the effectiveness of the main components (e.g., task experts and universal expert) in the proposed MoCLE method in Sec.
-4.4.1.
-Then in Sec.
-4.4.2 we conduct a thorough analysis to study how the proposed MoCLE method would respond to the change of hyper-parameters (e.g., temperature and the number of clusters and task experts).
+In this section, we first ablate the effectiveness of the main components (e.g., task experts and universal expert) in the proposed MoCLE method.
+Then we conduct a thorough analysis to study how the proposed MoCLE method would respond to the change of hyper-parameters (e.g., temperature and the number of clusters and task experts).
 Notice that for the ablations, we use Flickr30K (Flickr), Hateful Memes (HM), ScienceQA (SciQA), IconQA, Visual Spatial Reasoning (VSR) and TextVQA for evaluation.
-Components Flickr HM SciQA IconQA VSR TextVQA
-InstructBLIP 81.3 65.3 62.0 43.4 60.8 53.9
-+LoRA (rank=8) 81.3 65.1 57.4 44.2 62.8 49.4
-+LoRA (rank=64) 81.5 65.2 62.0 43.9 62.6 49.0
-+Mixture (Cluster) 81.9 65.4 63.3 46.1 58.9 54.9
-+Universal Expert 81.9 65.6 63.9 46.3 64.7 57.1
+Components Flickr HM SciQA IconQA VSR TextVQA Avg.
+InstructBLIP 81.3 65.3 62.0 43.4 60.8 53.9 61.1
++LoRA (rank=8) 81.3 65.1 57.4 44.2 62.8 49.4 60.0
+ (rank=64) 81.5 65.2 62.0 43.9 62.6 49.0 60.7
++Mixture (Cluster) 81.9 65.4 63.3 46.1 58.9 54.9 61.8
++Universal Expert 81.9 65.6 63.9 46.3 64.7 57.1 63.3
 Table 3: Comparison of individual components of the MoCLE framework in zero-shot vision-language tasks.
-We include three ablation studies that cumulatively add (i) parameter-efficient finetuning of LLM using LoRA of different ranks, (ii) replacing the single LoRA with its mixture and route inputs by their cluster, and (iii) training with the universal expert.
+We include three ablation studies that cumulatively add (i) parameter-efficient finetuning of LLM using LoRA of different ranks (higher ranks are ablated, i.e.
+rank=64, but we use rank=8 for ablations and MoCLE), (ii) replacing the single LoRA with its mixture and route inputs by their cluster, and (iii) training with the universal expert.
 Default settings are marked in gray.
-4.4.1 Effects of Different Components
-We start from vanilla InstructBLIP (Vicunna-7B) and add the main components in the proposed MoCLE method step by step.
-As shown in Table 2, naively applying LoRA to InstructBLIP could be problematic.
+4.4.1 Effects of Different Components.
+We start from vanilla InstructBLIP with Vicuna-7B and add the main components in the proposed MoCLE method step by step.
+As shown in Table 3, naïvely applying LoRA to InstructBLIP could be problematic.
 Although we observe minor performance gain in IconQA and VSR, the peformance on ScienceQA and TextVQA are much inferior to InstructBLIP.
@@
 However, the resulting model is still outperformed by InstructBLIP with less trainable parameters.
-This indicates that naively increasing the capacity of the network cannot mitigate the task conflicts in the training data.
+This indicates that naïvely increasing the capacity of the network cannot mitigate the task conflicts in the training data.
 However, when we replace a single LoRA with the Mixture of LoRAs and route the input based on the cluster of their instructions, the resulting model outperforms its predecessor by a large margin on SciQA, IconQA, and TextVQA.
@@
 This could be due to the fact that instruction-tuned model generalizes to unseen tasks by training on many instructions, while in our case, each expert sees fewer instructions than the original dense model does.
-As shown in Table 2, this issue could be alleviated by the use of the universal expert, which is updated by the gradients from all the training data.
+As shown in Table 3, this issue could be alleviated by the usage of the universal expert, which is updated by the gradients from all the training data.
 The same result can be observed in TextVQA, which requires not only the VQA ability but also additional optical character recognition (OCR) skills.
 Specifically, in the training stage, we have instructions “Question: {}.
-Short answer:” for all the VQA datasets and “OCR tokens: {}.
-Write a description for the photo” for TextCaps [49], but the instruction of TextVQA [50] is “OCR tokens: {}, Question: {}.
-Short answer:”.
-This requires the model to combine the knowledge from two tasks.
-Universal Expert vs. Top-2 Experts.
+Short answer:” for all VQA datasets and “OCR tokens: {}.
+Write a description for the photo” for TextCaps [55], while the TextVQA [56] instruction is “OCR tokens: {}, Question: {}.
+Short answer:”, requiring LVLMs to combine knowledge from both tasks.
+4.4.2 Universal Expert vs. Top-2 Experts.
 MoCLE shares similarities with the top-2 variant of MoE.
@@
 As shown in Table 4, the top-2 MoE model yields inferior results compared to MoCLE.
-Therefore the naively activating one more expert cannot address the generalization problem.
+Therefore the naïvely activating one more expert cannot address the generalization problem.
 Surprisingly, this variant performs even worse than the MoCLE model without the universal expert reported in Table 3.
 This might be resulted from task conflicts with more parameters shared.
-Flickr HM SciQA IconQA VSR TextVQA
-Universal 81.9 65.6 63.9 46.3 64.7 57.1
-top-2 82.0 64.7 61.9 45.5 56.3 52.0
+Flickr HM SciQA IconQA VSR TextVQA Avg.
+Universal 81.9 65.6 63.9 46.3 64.7 57.1 63.3
+Top-2 82.0 64.7 61.9 45.5 56.3 52.0 60.4
 Table 4: Ablation study on the universal expert by comparing with either (i) a universal expert that is activated all the time or (ii) expert with the second largest logit, in addition to the top-1 expert.
-Gating strategy Flickr HM SciQA IconQA VSR TextVQA
-Token 81.7 65.4 61.9 44.0 49.0 46.6
-Sentence 82.0 65.1 62.3 45.3 56.6 47.0
-Cluster 81.9 65.6 63.9 46.3 64.7 57.1
+Gating strategy Flickr HM SciQA IconQA VSR TextVQA Avg.
+Token 81.7 65.4 61.9 44.0 49.0 46.6 58.1
+Sentence 82.0 65.1 62.3 45.3 56.6 47.0 59.7
+Dataset 80.3 64.6 63.1 45.9 62.6 55.8 62.1
+Cluster 81.9 65.6 63.9 46.3 64.7 57.1 63.3
 Table 5: Ablation study on routing inputs based on different conditions. Default settings are marked in gray.
-Gating Strategies.
-We compare the proposed cluster-conditioned gating strategy with two popular ones: Token-MoLE and Sentence-MoLE, where MoLE denotes the mixture of LoRA experts by applying the MoE technique to LoRA, Token-MoLE obtains the routing decision based on the hidden representations of each token, and Sentence-MoLE averages the representations of the instruction tokens while excluding the visual tokens.
+4.4.3 Gating Strategies.
+We compare the proposed cluster-conditioned gating strategy with Token-MoLE, Sentence-MoLE and Dataset-MoLE, where MoLE denotes the mixture of LoRA experts by applying the MoE technique to LoRA, Token-MoLE obtains the routing decision based on the hidden representations of each token, and Sentence-MoLE averages the representations of the instruction tokens while excluding the visual tokens.
+Dataset-MoLE is a special case of MoCLE as it treats each dataset as a cluster while MoCLE leverages $k$-means to obtain clusters among all the datasets.
+Note that the universal expert is adopted for all these MoLEs.
 As observed in Table 5, Token-MoLE and Sentence-MoLE give inferior results on the evaluation tasks.
-We speculate this is because that (1) an expert only learns a subset of the tasks, giving poor generalizations and (2) similar tasks are not grouped together by the same expert, resulting in task conflicts within that expert.
+We speculate that this is because (1) an expert only learns a subset of the tasks, giving poor generalization and (2) similar tasks are not grouped together by the same expert, resulting in task conflicts within that expert.
 The later can be evidenced by the routing visualization in Figure 4, where samples in the same dataset are routed to multiple experts instead of a dedicate one.
-4.4.2 Hyper-parameter Analysis
-In this section, we analyze in details the effect of the hyper-parameter in the MoCLE framework to the performance.
-Specifically, we investigate the how would MoCLE react to the number of clusters as introduced in Sec.
-3.3 and the number of task experts as well as the temperature value in Sec.
-3.4.
-If not otherwise specified, we employ all the components ablated in Sec.
-4.4.1 and use a default configuration of 64 clusters with 4 experts and a temperature of 0.05.
+Further, we observe performance drops when replacing $k$-means clustering with statically treating each dataset as a cluster.
+This results from the fact that Dataset-MoLE is less flexible than MoCLE because it can only assign a dataset to one cluster.
+However, in practice, we observe multiple tasks in a dataset which should be assigned to different clusters.
+(e.g., llava_150k contains reasoning, conversations and captioning, which are assigned to different clusters/experts as in Figure 2(a) and 3(a)).
+Table 6: Ablation study on the number of clusters $K$.
 # Clusters Flickr HM SciQA IconQA VSR TextVQA
@@
 256 82.5 65.6 63.2 46.0 57.0 56.9
+Temperature $\tau$ Flickr HM SciQA IconQA VSR TextVQA
+0.01 81.6 64.7 61.7 46.3 61.2 55.8
+0.05 81.9 65.6 63.9 46.3 64.7 57.1
+0.1 82.3 65.7 63.7 45.7 63.0 56.7
+0.2 81.9 64.8 62.1 45.2 61.3 54.0
 Table 6: Ablation study on the number of clusters $K$.
+Table 7: Ablation study on the temperature $\tau$ in Eq. (1).
 #Task experts Flickr HM SciQA IconQA VSR TextVQA
@@
 Table 8: Ablation study on the number of task experts.
-Temperature $\tau$ Flickr HM SciQA IconQA VSR TextVQA
-0.01 81.6 64.7 61.7 46.3 61.2 55.8
-0.05 81.9 65.6 63.9 46.3 64.7 57.1
-0.1 82.3 65.7 63.7 45.7 63.0 56.7
-0.2 81.9 64.8 62.1 45.2 61.3 54.0
-Table 7: Ablation study on the temperature $\tau$ in Eq. (1).
-Number of Clusters.
+4.4.4 Number of Clusters.
 The merit of sentence encoding and clustering lies in reducing the complexity of the instruction, offering robust and semantic-relevant routing process.
@@
 So we use 64 clusters as the default setting.
-Number of Task Experts.
+4.4.5 Number of Task Experts.
 As demonstrated in Table 8, more task experts usually provides with stronger capacity.
@@
 Hence, we use 4 task experts as the default setting.
-Figure 3: t-SNE visualization of the instruction encoding. Different colors correspond to different datasets, while the shape of the markers indicates the task category defined manually.
-Temperature.
+4.4.6 Temperature.
 In the proposed MoCLE method, the value of temperature plays an important role in controlling the contribution of the universal expert.
@@
 4.5 Visualizations
-Figure 4: Clustering assignment of the training datasets when $K=64$.
-The labels on the vertical axis indicate the names of the datasets and they are arranged such that different datasets of the same task are placed together.
+Figure 3: (a) Clustering assignment of the training datasets when $K=64$.
+The labels on the vertical axis indicate the names of the datasets.
 Within each labeled block, multiple rows represent different subsets, each of which is associated with a unique task template.
 The horizontal axis denotes the cluster index to which these subsets are assigned.
-Darker color indicates more samples of this subset are assigned to certain cluster.
-Clustering.
+(b) t-SNE visualization of the instruction encoding.
+Different colors correspond to different datasets, while the shape of the markers indicates the task category defined manually.
+4.5.1 Clustering.
 We first show the justification to represent the training data via their instructions.
 Specifically, for each dataset, we sample 100 examples and encode their instructions with the all-MiniLM-L6-v2 variant of the Sentence Transformer model [49].
-After that, we visualize the data in Figure 3 via t-SNE [51].
-According to Figure 3, we observe that (1) samples from different datasets but from the same task are grouped together.
+After that, we visualize the data in Figure 2(b) via t-SNE [44].
+According to Figure 2(b), we observe that (1) samples from different datasets but from the same task are grouped together.
 For example, coco_vqg, aok_vqg, and ok_vqg are all VQG (visual question generation) tasks (triangle markers) and reside on the left part of the figure.
@@
 For example, coco_cap and textcaps both belong to the image captioning (CAP) task (small dots) and stay close to each other.
-Similarly, both the visual question answering (VQA, cross markers) task with datasets (e.g., aok_vqa, coco_vqa, ocrvqa, and ok_vqa) and conversation (CONV) task with datasets (e.g., llava_conversation, llava_detail, and llava_reason) involve answering user questions, which lie in the middle part of Figure 3, suggesting that instructions are good representatives of the training data.
+Similarly, both the visual question answering (VQA, cross markers) task with datasets (e.g., aok_vqa, coco_vqa, ocrvqa, and ok_vqa) and conversation (CONV) task with datasets (e.g., llava_conversation, llava_detail, and llava_reason) involve answering user questions, which lie in the middle part of Figure 2(b), suggesting that instructions are good representatives of the training data.
 We then cluster all the instructions of the examples in the training data into 64 groups using $k$-means clustering.
-Figure 4 shows the cluster assignment of the training data.
+Figure 2(a) shows the cluster assignment of the training data.
 Here, each row in the heatmap denotes a subset of a dataset.
@@
 Those evidences justify the use of clustering on task instructions as an automatic partition strategy for training datasets.
-Routing results.
-Figure 5 visualizes the routing decisions of the proposed MoCLE and Sentence-MoLE.
+4.5.2 Routing results.
+Figure 4 visualizes the routing decisions of our MoCLE and Sentence-MoLE.
 We obtain both results from one mixture of LoRA, i.e., one linear module in a layer.
-The routing results are aggregated by the subset of datasets similar to Figure 4.
+The routing results are aggregated by the subset of datasets similar to Figure 2(a).
 Shown in Figure 3(a), MoCLE could achieve task-level routing for the inputs.
@@
 Further, the remaining 3 experts are not dispensable as without them all tasks would clutter in a single expert, resulting in task conflicts.
-On the contrary, routing pattern of Sentence-MoLE in Figure 4(b) reveals little correlations between tasks/datasets and experts.
+Instead, routing pattern of Sentence-MoLE in Figure 3(b) reveals little correlations between tasks/datasets and experts.
 That is, different datasets/tasks obtain similar routing decisions, still suffering from task conflicts.
@@
 Figure 4: Routing decisions of one LoRA mixture for MoCLE and Sentence-MoLE.
-The setup of the vertical axis is similar to Fig.
-4 except that we also include the held-out tasks.
+The setup of the vertical axis is similar to Figure 2(a) except that we also include the held-out tasks.
 They are separated by a dotted line on the vertical axis.
@@
 In addition, we achieve task specialization and generalization in MoCLE simultaneously via a separate universal expert.
-Comprehensive evaluations of MoCLE on 12 unseen tasks show the effectiveness of the proposed MoCLE method.
-Acknowledgments.
-We gratefully acknowledge support of MindSpore, CANN (Compute Architecture for Neural Networks) and Ascend AI Processor used for this research.
+Comprehensive evaluations of MoCLE on 11 unseen tasks show the effectiveness of our MoCLE.
+Acknowledgement
+We gratefully acknowledge the support of MindSpore, CANN (Compute Architecture for Neural Networks) and Ascend AI Processor used for this research.
 Appendix
-Appendix A Sources of Results in Table 1
-For Flamingo-80B [2], Qwen-VL [3], MiniGPT-v2 [5], BLIP-2 [26], and InstructBLIP [11], the results are directly from their papers. For LLaVA and MiniGPT-4, we take the results from [5].
-Appendix B Performance on LVLM Benchmarks
-We evaluate MoCLE and the reproduced InstructBLIP on several LVLM benchmarks such as POPE [30], MME [13], and MM-Vet [59].
-In details, given an image, POPE evaluates the object hallucination of LVLM by determining whether it could correctly respond to the question such as “Is there a broccoli in the image?”.
-MME and MM-Vet investigate various capacities of LVLM such as perception, cognition, etc.
-As shown in Table 8, incorporating MoCLE into InstructBLIP significantly reduces the object hallucination.
-This might result from the fact that questions w.r.t various object categories belong to different instruction clusters [^1] More sample instructions in the clusters are in Sec.
-H.
-and are routed to appropriate task experts in MoCLE.
-In addition, MoCLE also improves various abilities in LVLM as demonstrated in the performance gains in MME and MM-Vet.
-Models POPE (adv.) POPE (rand.) POPE (pop.) MME (percep.) MM-Vet
-InstructBLIP 77.6 84.3 78.2 1202.9 23.6
-+ MoCLE 82.1 88.0 83.7 1222.6 26.0
-Table 8: Performance of MoCLE on LVLM benchmarks.
-We report the F1-score, scores in perceptions, and the overall score for POPE, MME and MM-Vet, where a larger value corresponds to the better performance.
-POPE (adv.) indicates the adversarial objects splits of POPE.
-Similarly, POPE (rand.) and POPE (pop.) are for random and popular splits, respectively.
-Appendix C Training Details
-Following [11], we adopt the same training configurations for the mentioned models such as the proposed MoCLE, the reproduced InstructBLIP, (i.e., InstructBLIP (${}^{{\ddagger}}$) and the task experts in Sec.
+Appendix 0.A Sources of Results in Table 1
+For LLaVA (13B) and MiniGPT-4 (13B), the results on GQA, VSR, OKVQA and IconQA are from the Table 3 of [5], the results on POPE and A-OKVQA (Direct) are from Table 3 and 6 of [33], respectively.
+In particular, the performance of MiniGPT-4 (13B) on MME is from its official leaderboard  [^1] https://github.com/BradyFU/Awesome-Multimodal-Large-Language-Models/tree/Evaluation.
+All the task performances of InstructBLIP (7B) are obtained from its reproduced version implemented in this paper.
+Appendix 0.B Training Details
+Following [14], we adopt the same training configurations for the mentioned models such as the proposed MoCLE, the reproduced InstructBLIP (7B) and the task experts in Sec.
 1.
@@
 We apply a linear warmup of the learning rate during the initial 1000 steps, increasing from ${10}^{-8}$ to ${10}^{-5}$, followed by a cosine decay with a minimum learning rate of 0.
-Appendix D Weights of the Universal Experts
-During training, if some training data obtain a very large weight on a task experts, such data tend to be very specific and might be less beneficial to other tasks.
+Appendix 0.C Weights of the Universal Experts
+During training, if some training data obtains a very large weight on a task expert, such data tend to be very specific and might be less beneficial to other tasks.
 Hence, they get less weight on the universal expert.
@@
 Figure 5: Weights of the universal expert for different datasets. Colors indicate different datasets.
-Appendix E Datasets for Vision-language Instruction Tuning
-Here we give a detailed description of the datasets used during instruction tuning.
+Appendix 0.D Datasets for Vision-language Instruction Tuning
+Here we give a detailed description of datasets used during instruction tuning.
 As shown in Table 9, we use 13 datasets in total.
-Here multiple datasets might be associated with the same data sources because these sources are formatted by different groups of task templates (Sec.
-F).
-For LLaVA-Instruct150K [33], we do not apply any task template as it is well formatted.
-For Web Capfilt, we use COCO+VG+CC+SBU (14M images) as introduced in [25].
-In regard to which split is used, please refer to the original InstructBLIP [11] for more details.
-During training, we sample the datasets using the data balancing strategy proposed in [11].
+Here multiple datasets might be associated with the same data sources because these sources are formatted by different groups of task templates (see Appendix 0.E).
+For LLaVA-Instruct150K [38], we do not apply any task template as it has been well formatted.
+For Web Capfilt, we use COCO+VG+CC+SBU (14M images) as [29].
+Regarding with which split is used, please refer to the InstructBLIP paper [14] for more details.
+During training, we sample datasets using the data balancing strategy as in [14].
+Appendix 0.E Task Templates
+Following [14], we use the same set of task templates for instruction tuning and held-in/out evaluation. Please refer to Tables 14 and 15 for training and evaluation templates.
+Appendix 0.F Case Studies
+In this section, we present several case studies with MoCLE.
+First, we study its conversation abilities via a range of tasks, including object counting, optical character recognition (OCR), and image introduction.
+Then we showcase some example instructions sampled from different clusters.
+0.F.1 Conversations
+In Table 10, we instruct the model to conduct a very difficult object counting task.
+The correct answer for this question is 63, which is quite hard for existing LVLMs.
+InstructBLIP fails to give the correct answers, while with MoCLE, InstructBLIP can respond the user query in a much more proper manner.
+In Table 11, the model is queried to recognize the character in the image.
+InstructBLIP performs not so well on this query possibly because OCR-related tasks conflict with other tasks during training.
+With MoCLE, the model can give correct results.
+In Table 12, we ask the model to introduce a famous person in the image.
+InstructBLIP gives a blunt response to the user query and does not follow the instruction of “introduction”.
+This might be due to the conflict between image caption and conversation tasks.
+In the training data, there are a large portion of image caption data that require the model to give a brief description to the image, while the user query in this example expects a detailed introduction to Albert Einstein.
+With MoCLE, the user query is identified and routed to the correct expert that is specialized at such a conversation task, thus, the model outputs a desired response.
+Similarly, in Table 13, we ask the model to describe the image in a detailed manner.
+InstructBLIP still mistakes this query as an image caption task and gives very short caption to this image.
+Instead, with MoCLE, the model correctly understands the “in details” in the instruction and provides sufficient details.
 Datasets Data Source Task Template Group
@@
 Table 9: List of datasets used for training.
-Appendix F Task Templates
-Following [11], we use the same set of task templates for instruction tuning and zero-shot evaluation. Please refer to Tables 14 and 15 for training and evaluation templates.
-Appendix G More Discussion
-Limitation.
-Although effective, we mainly focus on task conflicts among text-based conversation tasks (e.g., image captioning and VQA) in this paper, while the support of our MoCLE for more complicated visual perception tasks (e.g., object detection [17, 27]) is appealing, which has shown more severe task conflicts with the conversation tasks [68].
-Moreover, the applicability of MoCLE to image diffusion models [9, 14, 35, 29] is also a potential research direction, considering the ubiquity of negative transfer phenomenon.
 Model responses to counting-related queries.
@@
 InstructBLIP The exact characters in the image are “S 3000 CD.” +MoCLE The exact characters in the image are “S 3302 CD.” Table 11: Model responses to OCR-related queries.
-Appendix H Case Studies
-In this section, we present several case studies with MoCLE.
-First, we study its conversation abilities via a range of tasks, including object counting, optical character recognition (OCR), and image introduction.
-Then we showcase some example instructions sampled from different clusters.
-Conversations
-In Table 10, we instruct the model to conduct a very difficult object counting task.
-The correct answer for this question is 63, which is quite hard for existing LVLMs.
-InstructBLIP failed to give the correct answers, while with MoCLE, InstructBLIP can respond the user query in a much more proper manner.
-In Table G, the model is queried to recognize the character in the image.
-InstructBLIP performs not so well on this query possibly because OCR-related tasks conflict with other tasks during training.
-With MoCLE, the model can give correct results.
-In Table 12, we ask the model to introduce a famous person in the image.
-InstructBLIP gives a blunt response to the user query and does not follow the instruction of “introduction”.
-This might be due to the conflict between image caption and conversation tasks.
-In the training data, there are a large portion of image caption data that require the model to give a brief description to the image, while the user query in this example expects a detailed introduction to Albert Einstein.
-With MoCLE, the user query is identified and routed to the correct expert that is specialized at such a conversation task, thus, the model outputs a desired response.
-Similarly, in Table 13, we ask the model to describe the image in a detailed manner.
-InstructBLIP still mistakes this query as an image caption task and gives very short caption to this image.
-Differently, with MoCLE, the model correctly understands the “in details” in the instruction and provides sufficient details.
 Model responses when asked to introduce a famous person.
@@
 Table 12: Model responses to introduce a celebrity.
-Sample Instructions in Clusters
+0.F.2 Sample Instructions in Clusters
 In Table 16, we showcase some sample instructions assigned to different clusters.
-Though all the instructions in the 4 selected clusters belong to VQA-related tasks, they focus on various perspectives such as food, pet, men, and counting.
-This justifies the use of a large number of instruction clusters.
+Though all the instructions in the 4 selected clusters belong to VQA-related tasks, they focus on various perspectives such as food, pet, men, and counting, justifying the usage of a large number of instruction clusters.
 Moreover, in Table 7, we show that the model performs poorly with only a few clusters, i.e., $K=4$, and gets better as we increase $K$.
-With a few clusters, VQA, VQG, CONV and CAP tasks would occupy 4 clusters, leaving no more space for samples with different topics.
-As a result, the model might perform worse when mixing these topics.
+With a few clusters, VQA, VQG, CONV and CAP tasks will occupy 4 clusters, leaving no more space for samples with different topics.
+Thus, the model might perform worse when mixing these topics.
 Model responses when queried to give detailed image descriptions.
@@
 Table 13: Model responses to give detailed image description.
-Template Group Task Template
-CAP A short image caption:
-A short image description:
-A photo of
-An image that shows
-Write a short description for the image.
-Write a description for the photo.
-Provide a description of what is presented in the photo.
-Briefly describe the content of the image.
-Can you briefly explain what you see in the image?
-Could you use a few words to describe what you perceive in the photo?
-Please provide a short depiction of the picture.
-Using language, provide a short account of the image.
-Use a few words to illustrate what is happening in the picture.
-VQA {Question}
-Question: {Question}
-Question A short answer to the question is
-Q: {Question} A:
-Question: {Question} Short answer:
-Given the image, answer the following question with no more than three words. {Question}
-Based on the image, respond to this question with a short answer: {Question}. Answer:
-Use the provided image to answer the question: {Question} Provide your answer as short as possible:
-What is the answer to the following question? “{Question}”
-The question “{Question}” can be answered using the image. A short answer is
-VQG Given the image, generate a question whose answer is: {Answer}. Question:
-Based on the image, provide a question with the answer: {Answer}. Question:
-Given the visual representation, create a question for which the answer is “{Answer}”.
-From the image provided, craft a question that leads to the reply: {Answer}. Question:
-Considering the picture, come up with a question where the answer is: {Answer}.
-Taking the image into account, generate an question that has the answer: {Answer}. Question:
+Template Group
+Task Template
+CAP
+“A short image caption:”
+“A short image description:”
+“A photo of”
+“An image that shows”
+“Write a short description for the image.”
+“Write a description for the photo.”
+“Provide a description of what is presented in the photo.”
+“Briefly describe the content of the image.”
+“Can you briefly explain what you see in the image?”
+“Could you use a few words to describe what you perceive in the photo?”
+“Please provide a short depiction of the picture.”
+“Using language, provide a short account of the image.”
+“Use a few words to illustrate what is happening in the picture.”
+VQA
+“{Question}”
+“Question: {Question}”
+“Question A short answer to the question is”
+“Q: {Question} A:”
+“Question: {Question} Short answer:”
+“Given the image, answer the following question with no more than three words. {Question}”
+“Based on the image, respond to this question with a short answer: {Question}. Answer:”
+“Use the provided image to answer the question: {Question} Provide your answer as short as possible:”
+“What is the answer to the following question? {Question}”
+“The question {Question} can be answered using the image. A short answer is”
+VQG
+“Given the image, generate a question whose answer is: {Answer}. Question:”
+“Based on the image, provide a question with the answer: {Answer}. Question:”
+“Given the visual representation, create a question for which the answer is {Answer}.”
+“From the image provided, craft a question that leads to the reply: {Answer}. Question:”
+“Considering the picture, come up with a question where the answer is: {Answer}.”
+“Taking the image into account, generate an question that has the answer: {Answer}. Question:”
 […]
-Evaluation Dataset Instruction
-GQA, VizWiz, iVQA, MSVD, MSRVTT Question: {} Short answer:
-NoCaps, Flickr30k A short image description:
-TextVQA OCR tokens: {}. Question: {} Short answer:
-IconQA Question: {} Options: {}. Short answer:
-ScienceQA Context: {} Question: {} Options: {}. Answer:
-HatefulMemes This is an image with: “{}” written on it. Is it hateful? Answer:
-VSR Based on the image, is this statement true or false? “{}” Answer:
+Evaluation Dataset
+Instruction
+GQA, iVQA,
+“Question: {} Short answer:”
+MSVD, MSRVTT,
+OKVQA, VQAV2,
+A-OKVQA(Direct)
+A-OKVQA(MC),
+“Question: {} Options: {}. Short answer:”
+IconQA
+NoCaps, Flickr30k
+“A short image description:”
+TextVQA
+“OCR tokens: {}. Question: {} Short answer:”
+ScienceQA
+“Context: {} Question: {} Options: {}. Answer:”
+HatefulMemes
+“This is an image with: “{}” written on it. Is it hateful? Answer:”
+VSR
+“Based on the image, is this statement true or false? “{}” Answer:”
+Visdual Dialog
 […]
 Table 15: Task templates used during evaluation.
-Cluster Instruction Samples Topics
-1 “Q: what is being done to the food in the glass fronted box? A:” Food
+Cluster
+Instruction Samples
+Topics
+1
+“Q: what is being done to the food in the glass fronted box? A:”
+Food
 “Q: what category of pizza would this fall into? Options: (a) vegetarian (b) meat lovers (c) pesto (d) pepperoni, A:”
@@
 “what category of pizzas would this be considered?”
-2 “Q: what sport is the cartoon dog playing? A:” Pet
+2
+“Q: what sport is the cartoon dog playing? A:”
+Pet
 “Question: what is likely her favorite animal? Options: (a) cat (b) dog (c) pig (d) sheep, Short answer:”
@@
 “What might the relationship between the two women and the dog be?”
-3 “what type dressing does this man favor?” Men
+3
+“what type dressing does this man favor?”
+Men
 “Based on the image, respond to this question with a short answer: what are the men doing?. Answer:”
@@
 “Question: why is the man kneeling on the ground?”
-4 “how many more animals need to be added to all of these to get the number ten?” Counting
+4
+“how many more animals need to be added to all of these to get the number ten?”
+Counting
 “Question: how many big elephants are inside of this zoo enclosure together? Options: (a) one (b) four (c) two (d) three, Short answer:”
@@
 Table 16: Sampled instructions from different clusters.
+Appendix 0.G Limitation
+Although effective, we mainly focus on task conflicts among text-based conversation tasks (e.g., image captioning and VQA [18]) in this paper, while the support of our MoCLE for more complicated visual perception tasks (e.g., object detection [20, 30]) is appealing, which has shown more severe task conflicts with the conversation tasks [74].
+Moreover, the applicability of MoCLE to image diffusion models [9, 17, 39, 32] is also a potential research direction, considering the ubiquity of negative transfer phenomenon.

### v5

@@
-Project Page: https://gyhdog99.github.io/projects/mocle/
+Mixture of Cluster-Conditional LoRA Experts for Vision-Language Instruction Tuning
 Abstract
@@
 A separate universal expert is further incorporated to improve generalization capabilities of MoCLE for novel instructions.
-Extensive experiments on 11 zero-shot tasks demonstrate the effectiveness of MoCLE.
-Keywords:
-Instruction tuning Mixture of Experts Vision and language
+Extensive experiments on InstructBLIP and LLaVA demonstrate the effectiveness of MoCLE.
+Mixture of Cluster-Conditional LoRA Experts
+for Vision-Language Instruction Tuning
+Yunhao Gou1,2*, Zhili Liu2,3*, Kai Chen2*, Lanqing Hong3, Hang Xu3, Aoxue Li3, Xin Jiang3, Zhenguo Li 3, Dit-Yan Yeung2, James T.
+Kwok2, Yu Zhang1,4, 1Southern University of Science and Technology, 2The Hong Kong University of Science and Technology, 3Huawei Noah’s Ark Lab, 4Peng Cheng Laboratory, Equal contribution.
+Correspondence: yu.zhang.ust@gmail.com
 1 Introduction
-There has been a continuously increasing trend to develop intelligent assistants that can follow human instructions [4, 48, 8], with instruction tuning emerging as a notably effective approach.
-This method leverages large-scale well-formatted instruction data to empower Large Language Models (LLMs) to execute various human instructions, showcasing their ability to generalize across novel unseen tasks [41].
-Likewise, efforts have been made to introduce similar capabilities to Large Vision-language Models (LVLMs) [3, 66, 63, 10, 5], including LLaVA series [38, 37], MiniGPT-4 [73] and InstructBLIP [14].
-They connect a trainable visual encoder with a frozen LLM and endow the LVLMs with the instruction-following ability by finetuning on diverse multi-modal instruction datasets.
-It is observed in both the LLMs [51, 59, 13] and LVLMs [3, 69, 31] that the ability to generalize to novel unseen instructions necessitates multi-task instruction tuning, i.e., training on a diverse collection of instruction-following tasks.
-However, the complexity of various instruction tasks brings difficulties for the model fine-tuning.
-Specifically, [60] find that under certain model sizes, multi-task instruction tuning even fails to bring performance gains for zero-shot tasks compared to the original models, mainly attributed to the negative transfer phenomenon [68] during multi-task instruction tuning, where the model struggles to optimize the losses of multiple conflicted tasks, leading towards the sub-optimal experimental performance for these tasks.
+There has been a continuously increasing trend to develop intelligent assistants that can follow human instructions Brown et al.
+(2020); OpenAI (2022); Chen et al.
+(2023c), with instruction tuning emerging as a notably effective approach.
+This method leverages large-scale well-formatted instruction data to empower Large Language Models (LLMs) to execute various human instructions, showcasing their ability to generalize across novel unseen tasks Longpre et al.
+(2023).
+Likewise, efforts have been made to introduce similar capabilities to Large Vision-language Models (LVLMs) Bai et al.
+(2023); Zhang et al.
+(2023); Ye et al.
+(2023); Chen et al.
+(2023e, a), including LLaVA series Liu et al.
+(2023b, a), MiniGPT-4 Zhu et al.
+(2023) and InstructBLIP Dai et al.
+(2023).
+It is observed that for both the LLMs Sanh et al.
+(2021); Wang et al.
+(2022); Chung et al.
+(2022) and LVLMs Bai et al.
+(2023); Zhao et al.
+(2023); Li et al.
+(2023b), the ability to generalize to novel unseen instructions necessitates multi-task instruction tuning, i.e., training on a diverse collection of instruction-following tasks.
+However, the complexity of various instruction tasks brings difficulties for model fine-tuning.
+Specifically, Wei et al.
+(2021) find that for certain model sizes, multi-task instruction tuning even fails to bring performance gains for zero-shot tasks compared to the original models.
+This is mainly attributed to the negative transfer phenomenon Zhang and Yang (2017) during multi-task instruction tuning, where the model struggles to optimize the losses of multiple conflicted tasks, leading to sub-optimal performance.
 Figure 1: Performance of the instruction-finetuned LVLMs on zero-shot tasks, where larger values indicate better performance.
 Only 2 out of 7 tasks benefit from instruction tuning from all the data, while the task experts show better performance on the other 5 tasks (i.e., Flickr 30K, GQA, HM, SciQA and IconQA).
-Similarly, tasks for vision-language instruction tuning (e.g., visual question answering and image captioning) focus on different perspectives of LVLMs, and thus, resulting in conflicts when sharing all the parameters as most studies adopt.
-In our preliminary study, we first split the instruction data into two disjoint subsets, including “cap” for image captioning and “vqa” for visual question answering.
-We then train InstructBLIP [14] using LoRA [21] on 3 kinds of data: “cap”, “vqa” and “full” (the full data) separately, to obtain 3 set of parameters (i.e., task experts).
-Following the held-out evaluation protocol [14], we evaluate these experts on the unseen datasets/tasks with the best expert.
+Similarly, tasks for vision-language instruction tuning (e.g., visual question answering and image captioning) focus on different perspectives of LVLMs.
+This results in conflicts as most studies adopt sharing of all parameters.
+In our preliminary study, we split the instruction data into two disjoint subsets (“cap” for image captioning, and “vqa” for visual question answering).
+We then train InstructBLIP Dai et al.
+(2023) using LoRA Hu et al.
+(2021) on three data sets (“cap”, “vqa” and the full data “full”) to obtain three sets of parameters (i.e., task experts).
+Following the held-out evaluation protocol Dai et al.
+(2023), we evaluate these experts on the unseen datasets/tasks with the best expert.
 As shown in Figure 1, on 5 out of the 7 downstream tasks, the InstructBLIP instruction-tuned on all the data is outperformed by the task expert finetuned with only a subset of data.
-Among the 5 tasks, the Flickr30k belongs to “cap”, and SciQA, GQA and IconQA belong to “vqa”, revealing that instruction tuning on the similar tasks brings positive transfer to downstream tasks, while training on the full data with dissimilar tasks could hurt the generalization performance.
-The disjoint task expert above is a naïve solution to negative transfer, where we manually partition training tasks and train each expert individually.
-However, there are several limitations of this strategy: (1) The taxonomies such as “vqa” and “cap” require human expertise and become difficult to scale as the number of tasks grows.
+Among the 5 tasks, Flickr30k belongs to “cap”, and SciQA, GQA and IconQA belong to “vqa”.
+This shows that instruction tuning on similar tasks brings positive transfer to downstream tasks, while training on the full data with dissimilar tasks can hurt generalization performance.
+The use of disjoint task experts above is a naïve solution to negative transfer, where we manually partition the training tasks and train each expert separately.
+However, it has several limitations: (1) The taxonomies such as “vqa” and “cap” require human expertise, and are difficult to scale as the number of tasks grows.
 (2) The ability to generalize to unseen tasks is inhibited, as we do not know which expert to choose for novel tasks, while some new tasks might benefit from multiple training tasks (e.g., VSR and TextVQA as in Figure 1).
 In this regard, specialization and generalization of LVLMs becomes a dilemma.
-Our work aims to develop an automatic and practical partition strategy and a network architecture that can strike a balance between specialization and generalization.
-We propose Mixture of Cluster-conditional LoRA Experts (MoCLE) for vision-language instruction tuning.
-Under the proposed MoCLE framework, we first cluster instructions of all the training data into several centroids via a pre-trained clustering model.
-In this way, similar tasks that could bring positive transfer to each other are automatically grouped into the same cluster, and different tasks that may cause conflict are separated (more justification for the usage of instruction clusters is detailed in Sec.
-3.3).
-Then we construct several task experts with each concentrating on the specific cluster.
-Using the cluster as a condition, a router dispatches the input data to one of the specialized task experts and a universal expert that is shared among all the data.
-As we activate a specialized expert for a group of similar tasks, tasks that are less alike are learned via separate experts, mitigating task conflicts.
+This paper aims to develop an automatic and practical partition strategy and a network architecture that strikes a balance between specialization and generalization.
+In particular, we propose the Mixture of Cluster-conditional LoRA Experts (MoCLE) for vision-language instruction tuning.
+In this proposed framework, we first cluster instructions of all the training data into several clusters via a pre-trained clustering model.
+In this way, similar tasks that can bring positive transfer to each other are automatically grouped into the same cluster, while different tasks that may cause conflict are separated (more justifications for the use of instruction clusters are detailed in Sec.
+3.2).
+Then we construct several task experts, with each focusing on a specific cluster.
+Using the cluster as condition, a router dispatches the input data to one of the specialized task experts and an universal expert that is shared among all data.
+As we activate a specialized expert for a group of similar tasks, tasks that are less similar are learned via separate experts, mitigating task conflicts.
 Meanwhile, since the universal expert trained on all tasks also contributes to the model outputs, we can enjoy generalization and specialization simultaneously.
-We validate effectiveness of MoCLE on the state-of-the-art InstructBLIP [14].
-MoCLE achieves significant improvement on the held-out tasks [14] including the image captioning (e.g., Flickr30K [64]), the text-based visual question answering (e.g., TextVQA [56]), the knowledge-based visual question answering (e.g., ScienceQA [42]), and novel visual reasoning (e.g., Visual Spatial Reasoning [36]).
-The main contributions of this work contain three parts.
+We validate effectiveness of MoCLE on InstructBLIP Dai et al.
+(2023) and LLaVA-1.5 Liu et al.
+(2023a) and observe remarkable performance gains compared to dense models and other MoE baselinesChen et al.
+(2023f, 2024).
+The main contributions of this work contain the following three parts,
   1. 1.
-    We identify the negative transfer phenomenon[40, 71] of task conflicts during the instruction tuning of LVLMs.
+    We identify the negative transfer phenomenon Liu et al. (2022b); Zhili et al. (2023) as tasks conflict during instruction tuning of LVLMs.
   2. 2.
-    We propose Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel parameter-efficient finetuning framework suitable for the vision-language instruction tuning, to mitigate the task conflicts and enjoy the benefits of huge data training simultaneously.
+    We propose Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel parameter-efficient finetuning framework suitable for the vision-language instruction tuning, to mitigate task conflicts and enjoy the benefits of huge data training simultaneously.
   3. 3.
-    Our proposed MoCLE achieves remarkable performance gains on the unseen held-out tasks compared with a strong baseline model InstructBLIP [14].
+    Our proposed MoCLE achieves remarkable performance gains on held-in/out tasks compared to dense models and other MoE baselinesChen et al. (2023f, 2024).
 2 Related Work
-2.0.1 Multi-task Instruction Tuning.
-Instruction tuning, as elucidated in [51, 60], is a technique where a language model is finetuned over a collection of tasks and instruction templates that tells the model the intention of the task.
-The primary goal of this technique is to teach the model to understand the relationship between the instruction and the input/output pairs, so that the model could generalize to unseen task given novel instructions.
-Increasing the number of diverse instructions [51], the number of tasks [59, 13], and diversity of data [72] have been shown to improve the performance.
-However, [60] observe under certain model sizes, instruction tuning of language models fail to bring performance gains to unseen tasks when compared with untuned base models.
-They argue this is due to the fact all model capacity is used to learn the mixture of instruction tuning tasks.
-In this work, we aim to mitigate such issue in vision-language instruction tuning by the specialized experts.
-2.0.2 Large Multi-modal Models.
-To inherit strong reasoning and generalization abilities of LLMs and meanwhile understand visual inputs, recent works including Flamingo [2], BLIP-2 [28], LLaMA-adapter [67] connect LLMs to the visual modality with a frozen visual encoder and a trainable module, which is then finetuned with plain image-text pairs.
-To further improve instruction following abilities and response quality, LLaVA series [38, 37] finetune the whole LLM on the vision-language conversational data generated by GPT-4.
-MiniGPT-4 [73] performs training on ChatGPT-generated image captions instead of traditional caption data that are shorter and informational sparser.
-InstructBLIP [14] collects a wide range of academic VQA datasets and manually crafts 10 to 15 distinct instruction templates for each kind of task.
-In this paper, we primarily consider InstructBLIP as the base model.
-2.0.3 Mixture of Experts.
-MoE models [23, 24, 53] are considered as an effective way to increase the model capacity in terms of parameter size.
-Within MoEs, certain parts of the model are activated while the computation is kept the same or close to its dense counterpart.
-Recently, it has been thoroughly investigated in the field of computer vision[50, 71], natural language processing[27, 15] and multi-modal learning[47, 54, 12, 11, 34].
-MoCE [71] tackles negative transfer in the self-supervised visual pre-training[6, 7] using a mixture of cluster experts.
-Different from MoCE requiring to be finetuned on downstream tasks, our MoCLE studies the zero-shot generalization to unseen tasks.
-Another line of works combines the MoE with adapters.
-AdaMix[58] randomly selects adapters during training and adopts the average of expert parameters during inference.
-Zadouri et al.
-[65] propose mixture of LoRAs for instruction tuning of LLMs.
-Concurrently, MoE-LLaVA [34] sparsifies LVLMs via MoEs for more efficient model scaling.
-The most relevant to ours include the Octavius [12] and LLaVA-MoLE [11], which also address task conflicts in vision-language instruction tuning by applying the MoE architecture with LoRAs.
-However, they do not explore the issue of instruction generalization in the architecture designs.
-Instead, we design a novel Mixture of Cluster-conditional LoRA Experts together with the universal experts, which is the first work addressing both task conflicts and generalization with a unified architecture.
+2.1 Multi-Task Instruction Tuning
+Instruction tuning Sanh et al.
+(2021); Wei et al.
+(2021) fine-tunes a language model across tasks and instruction templates to convey task intentions.
+Its goal is to teach the model to understand relationships between instructions and input/output pairs, enabling generalization to unseen tasks with novel instructions.
+Increasing the number of instructions Sanh et al.
+(2021), tasks Wang et al.
+(2022); Chung et al.
+(2022), and data diversity Zhou et al.
+(2023) have shown to be effective in improving performance.
+However, Wei et al.
+(2021) find that for certain model sizes, instruction tuning fails to outperform untuned models on unseen tasks due to full capacity utilization for learning task mixtures.
+Our work addresses this in vision-language instruction tuning using specialized experts.
+Figure 2: Overall pipeline of MoCLE.
+2.2 Mixture of Experts (MoE)
+MoE models Jacobs et al.
+(1991); Jordan and Jacobs (1993); Shazeer et al.
+(2017) are renowned for their ability to increase model capacity through parameter expansion.
+Recent research integrates MoE with adapters, exploring how pretrained adapters can be effectively combined Wu et al.
+(2024b), and how they enhance performance in both few-shot Huang et al.
+(2023) and zero-shot scenarios Jang et al.
+(2023); Muqeeth et al.
+(2024).
+Another line of research Chen et al.
+(2024); Wu et al.
+(2024a); Luo et al.
+(2024); Zadouri et al.
+(2023); Chen et al.
+(2023f) focuses on augmenting model capacity in a parameter-efficient manner.
+However, these approaches do not explicitly incorporate task/domain priors during the routing process, which might be limited when handling task conflicts.
+Two concurrent studies Liu et al.
+(2024); Li et al.
+(2024) incorporate domain information for expert routing.
+Unlike our approach, however, they do not address zero-shot generalization to unseen tasks.
 3 Methodology
-In this section, we start with the formulation of LVLM instruction tuning and the analysis of the limitations of task experts.
-We then introduce the proposed MoCLE in detail, with the overall framework shown in Figure 2.
-Figure 2: Overall pipeline of MoCLE.
-The left part is to cluster training data (a collection of vision-language tasks with diverse instructions) into several groups based on the instructions via $k$-means clustering algorithm.
-On the right is the MoCLE framework.
-$\mathbf{x}$ is the hidden representation of the sample (e.g., a sequence of tokens for images, instructions and the raw text inputs in the LLM) which is handled by 3 modules, including the universal expert, task expert, and the original linear module of the LLM.
-After that, the outputs of these modules are aggregated and passed to the next layer.
-Only experts, the cluster embedding and the gate parameters are updated during training, while parameters of the LLM are frozen.
-We insert such structure into the query and value projection in self-attention blocks of each layer by default.
+In this section, we start with the formulation of LVLM instruction tuning and an analysis of the limitations of task experts.
+We then introduce the proposed MoCLE.
+The overall framework is shown in Figure 2.
 3.1 Problem Formulation
-Suppose that there is a set of datasets that are divided into held-in and held-out datasets [14], where a large vision-language model is first finetuned on the held-in and then evaluated on the held-out datasets in a zero-shot manner.
-To unify and diversify input-output formats and promotes in instruction tuning, several task templates $\{T_{i}\}$ are designed to wrap the raw inputs, a pair of text and image $(X_{\text{txt}},X_{\text{img}})$ from the corresponding dataset, where $T_{i}$ denotes the $i$-th template.
+Suppose that there is a set of datasets that are divided into held-in and held-out datasets Dai et al.
+(2023).
+A large vision-language model is first finetuned on the held-in dataset, and then evaluated on the held-out dataset in a zero-shot manner.
+To unify and diversify input-output formats and promotes instruction tuning, several task templates $\{T_{i}\}$ are designed to wrap the raw inputs, which is a pair of text $X_{\text{txt}}$ and image $X_{\text{img}}$ from the dataset.
 For example, “Given the image, answer the question with no more than three words.
@@
 The instruction is defined as $I\equiv T_{i}(X_{\text{txt}})$ that wraps text inputs using the template.
-3.2 Limitation of Task Experts
-Figure 1 indicates that allocating task experts based on the categories of training datasets brings positive transfer to similar downstream tasks.
-Nevertheless, this approach has several notable limitations.
-First, taxonomies such as “vqa” and “cap” require delicate human designs with expertise in the task performed.
-Another limitation is the challenge of selecting the appropriate expert for a newly introduced task when testing.
-Further, with tasks trained in an isolated manner, knowledge sharing among different tasks is prevented, compromising generalization to unseen tasks and instructions as some might benefit from multiple tasks.
-Hence, we propose MoCLE in the following sections to address these limitations.
 3.2 Clustering Data by Instructions
 The purposes of partitioning the training data are two-fold.
-We target at training a task expert within a collection of similar tasks so as to prevent task conflicts.
-Meanwhile, we expect novel tasks to be assigned to the proper experts based on their cluster without manual intervention.
-To achieve these, we should find a proper description for the task.
-Instructions serve as the foundation for LVLMs because they encapsulate the commands and intentions that guide responses of the model.
-The instructions $I$ in LVLMs consist of templates and the raw inputs.
-By understanding the templates in the instructions, the model can discern the type of task it needs to perform.
-The raw inputs further infuse this framework with the substantive content that the model must process and respond to.
-Therefore, by conducting clustering on the instructions, we can partition training data without human labor into several clusters, each containing a set of similar tasks, which is further verified through experiments in Sec.
-4.5.
-Moreover, even when novel instruction tasks come, we can obtain a cluster index for them using the pre-trained clustering model without human intervention.
-Formally, denoting $\mathcal{E}(\cdot)$ as a pre-trained sentence encoder, $\mathbf{e}_{i}=\mathcal{E}(I_{i})$ is the sentence representation of an instruction $I_{i}$, we use the $k$-means clustering algorithm to group all instructions in the training datasets into $K$ clusters by iteratively minimizing $\sum_{j=1}^{K}\sum_{\mathbf{e}_{i}\in S_{j}}\|\mathbf{e}_{i}-\mathbf{c}_{j}\|^{2}$, where $S_{j}$ denotes the set of instructions assigned to the $j$-th cluster and $\mathbf{c}_{j}$ denotes the centroid of the $j$-th cluster.
-In each iteration of the $k$-means clustering algorithm, each instruction is assigned to the nearest centroid with all centroids updated as the average of the instruction representations in the corresponding cluster.
+First, we hope to train a task expert with a collection of similar tasks so as to avoid task conflicts.
+Second, we expect novel tasks to be automatically assigned to the proper experts based on their cluster without manual intervention.
+To achieve these goals, we conduct clustering on the instructions as they serve as the foundation for identifying different tasks.
+Formally, let $\mathcal{E}(\cdot)$ be a pre-trained sentence encoder, and $\mathbf{e}_{i}=\mathcal{E}(I_{i})$ be the sentence representation of an instruction $I_{i}$.
+We use the $k$-means clustering algorithm to group all instructions in the training datasets into $K$ clusters by iteratively minimizing $\sum_{j=1}^{K}\sum_{\mathbf{e}_{i}\in S_{j}}\|\mathbf{e}_{i}-\mathbf{c}_{j}\|^{2}$, where $S_{j}$ is the set of instructions assigned to the $j$th cluster, and $\mathbf{c}_{j}$ is the centroid of the $j$th cluster.
+In each $k$-means clustering iteration, each instruction is assigned to the nearest centroid with all centroids updated as the average of instruction representations in the corresponding cluster.
 3.3 Mixture of Cluster-Conditional LoRA Experts
 In addition to considerations at the data level, we also suggest an architectural design to tackle the issue of negative transfer.
-We propose Mixture of Cluster-conditional LoRA Experts (MoCLE) that learns to activate the LoRA expert at each layer given the cluster of the data.
-Specifically, denoting $E$ as the number of experts, we introduce a gate vector $\mathbf{G}\in\mathbb{R}^{E}$ that, given any input $\mathbf{x}_{i}$, determines the experts to which the input is routed.
-The gate vector is calculated as,
+We propose the Mixture of Cluster-conditional LoRA Experts (MoCLE) that learns to activate the LoRA expert at each layer given the cluster of the data.
+Specifically, let $E$ as be the number of experts.
+We introduce a gate vector $\mathbf{G}\in\mathbb{R}^{E}$.
+Given an input $\mathbf{x}_{i}$, $\mathbf{G}$ determines the experts to which the input is routed.
+The gate vector is obtained as:
 $\mathbf{G}=\text{top}_{k}\left(\text{softmax}\left(\frac{1}{\tau}\left(\mathbf{W}_{\text{gate}}\mathbf{C}_{[\mathbf{x}_{i}]}+\boldsymbol{\epsilon}\right)\right)\right),$ (1)
-where $\text{top}_{k}(\cdot)$ keeps the $k$ largest entries unchanged and sets others to zero and $\mathbf{C}_{[\mathbf{x}]}$, which is shared among all the layers, denotes the learnable embedding of the cluster that $\boldsymbol{x}$ belongs to.
+where $\text{top}_{k}(\cdot)$ keeps the $k$ largest entries unchanged and sets the others to zero.
+$\mathbf{C}_{[\mathbf{x}]}$, which is shared among all layers, is the learnable embedding of the cluster that $\mathbf{x}$ belongs to.
 This is the key for the model to choose proper task experts for the input data.
 To endow the clustering embedding with task information, we initialize it to be the centroid of the corresponding cluster.
-$\mathbf{W}_{\text{gate}}$ denotes trainable weights of the linear gate, $\boldsymbol{\epsilon}\sim N(0,\frac{1}{E})$ denotes a noise term that is designed to add some randomness to the expert choosing process, which encourages MoCLE to explore multiple combinations of experts during training.
-$\tau$ denotes a temperature hyperparamter.
-The output $\mathbf{y}_{i}$ is then computed as a sum of the weighted output of each expert and the original LLM linear layer [21] on the input $\mathbf{x}_{i}$ as,
+Moreover, $\mathbf{W}_{\text{gate}}$ is the trainable weights of the linear gate, which is learned at each layer where the MoE block is inserted, $\boldsymbol{\epsilon}\sim N(0,\frac{1}{E})$ is a noise term that adds randomness to the expert choosing process (and encourages MoCLE to explore multiple combinations of experts during training [^1] We do not apply load balancing during training as we found it might distort task specialization.), and $\tau$ is a temperature hyperparameter.
+The output $\mathbf{y}_{i}$ is then computed as the sum of weighted outputs of the experts, and the original LLM linear layer Hu et al.
+(2021) on the input $\mathbf{x}_{i}$, as:
 $\mathbf{y}_{i}=\sum_{e=1}^{E}G_{e}\mathbf{W}_{e}\mathbf{x}_{i}+\mathbf{W}_{0}\mathbf{x}_{i},$ (2)
-where $\mathbf{W}_{0}$ denotes the pre-trained linear layer of LVLMs, $\mathbf{W}_{e}$ is the linear projection weights of the $e$-th LoRA expert, and $G_{e}$, the $e$-th entry in $\mathbf{G}$, indicates the contribution of the $e$-th expert.
-In this formulation, the model learns to choose the expert that is suitable for the data by examining its cluster.
-We will show in Sec.
-4.4.3 that the condition of clusters in expert routing is indispensable for task generalization with significant superiority over typical routing conditions including sentence embeddings and token representations.
+where $\mathbf{W}_{0}$ is the pre-trained linear layer of LVLM, $\mathbf{W}_{e}$ is the linear projection weight of the $e$th LoRA expert, and $G_{e}$ (the $e$th entry in $\mathbf{G}$) indicates the contribution of the $e$th expert.
+Models LLM Expert Params.
+# Experts # Clusters Rank Temperature Trainable Params.
+InstructBLIP Vicuna-7B q_proj, v_proj 4 + 1(universal) 64 8 0.05 Q-Former, LoRAs LLaVA-1.5 Vicuna-1.5-7B up_proj, down_proj 4 + 1(universal) 4 128 0.1 MLP connector, LoRAs
+Table 1: Architecture details of MoCLE on different LVLMs. Note that experts are added to each layer of the transformer.
 3.4 Universal Expert
-However, as shown in Sec.
-4.4.1, we find that the formulation in Sec.
-3.4 still hurts the generalization ability of the entire model.
-We hypothesize that this might be attributed to the fact that the instruction-tuned models generalize to unseen tasks by training on extensive instructions [60], while in our formulation, each expert sees fewer instructions than the original dense model.
-Hence, we propose a universal expert that could learn from all the training data.
-Specifically, we fix the number of activated experts to 1 (i.e., $k$ equals 1) and define $G_{\text{max}}$ as the maximum element in $\mathbf{G}$.
-Then the output for all the experts is expressed as,
+As will be shown in Sec.
+4.4.1, the formulation in Sec.
+3.3 still hurts the generalization ability of the entire model, due to the fact that instruction-tuned models generalize to unseen tasks via training on extensive instructions Wei et al.
+(2021), while in our formulation, each expert sees fewer instructions than the original dense model.
+To alleviate this problem, we propose an universal expert that learns from all training data.
+Specifically, we fix the number of activated experts to 1 (i.e., $k$ in Eq.
+1 equals 1) and define $G_{\text{max}}$ as the maximum element in $\mathbf{G}$.
+Then the output for all the experts is expressed as:
 $\mathbf{y}_{i}=\left(\sum_{e=1}^{E}G_{e}\mathbf{W}_{e}+(1-G_{\text{max}})\mathbf{W}_{u}\right)\mathbf{x}_{i}+\mathbf{W}_{0}\mathbf{x}_{i}.$ (3)
-In this formulation, we additionally train a universal expert parameterized by $\mathbf{W}_{u}$.
-Different from task experts that are activated only for specific model inputs, the universal expert is activated for any inputs.
-The final output is the weighted sum of outputs from one of the experts and the universal expert puls the original LVLM’s output.
-The weighting for this universal expert is the complement of that of the chosen task expert.
+in which we additionally train an universal expert parameterized by $\mathbf{W}_{u}$.
+Different from the task experts that are activated only for specific model inputs, the universal expert is activated for all inputs.
+The final output is a weighted sum of outputs from one of the experts and the universal expert plus the original LVLM’s output.
 Consequently, the task expert learns distinct skills for certain tasks while the universal expert masters holistic understanding of the training corpus.
 The synergy between them offers both specialization and generalization for the LVLMs with MoCLE.
-4 Experiment
-In this section, we conduct an assessment of MoCLE across multiple downstream tasks under a zero-shot setting.
-Initially, we detail the experimental settings and implementation details, which is followed by a description of the datasets and instructions employed, along with the outcomes of our evaluations.
-Lastly, we present an ablation study and visualizations of clustering and routing results.
-4.1 Settings and Implementation Details
-In all the experiments, we utilize the InstructBLIP[14] model, particularly its Vicuna-7B variant [70], as the base LVLM to assess the effectiveness of MoCLE.
-We encode all the instructions of different datasets using the all-MiniLM-L6-v2 variant of the Sentence Transformer model [49] and cluster their embeddings into $64$ clusters via $k$-means clustering algorithm.
-We apply the proposed MoCLE on the LLM (i.e., Vicuna-7B) of InstructBLIP.
-Specifically, for each transformer layer in the LLM, we insert the MoCLE, which consists of 4 task LoRA experts and one LoRA universal expert, into the query and value projection of self-attention blocks.
-Inside each expert, we set the rank of the LoRA to 8 and the temperature $\tau$ to $0.05$.
-During instruction tuning, we only finetune parameters of Q-Former and the LoRA experts while keeping other parts of the LLM frozen.
-Models GQA VSR IconQA Visdial MME POPE A-OKVQA OKVQA VQAv2 Direct MC (test) (test-dev) LLaVA (13B) 41.3 51.2 43.0 - - 67.0 50.5 - 54.4 - MiniGPT-4 (13B) 30.8 41.6 37.6 - 581.7 71.4 38.7 - 37.5 - InstructBLIP (7B) 48.6 60.8 43.4 46.3 1202.9 77.6 58.8 73.8 57.0 77.4 49.3 64.7 46.3 46.9 1222.6 82.1 61.5 78.2 59.8 78.9 MoCLE (7B) (+0.7) (+3.9) (+2.9) (+0.6) (+19.7) (+4.5) (+2.7) (+4.4) (+2.8) (+1.5)
-Table 1: Zero-shot results on the held-out datasets (GQA, VSR, IconQA, Visdial) and LVLM benchmarks (MME, POPE) and evaluation on held-in datasets (A-OKVQA, OKVQA, VQAv2).
+Models GQA VSR IQA Visdial MME POPE A-OKVQA OKVQA VQAv2 Direct MC (test) (test-dev) InstructBLIP (7B) 48.6 60.8 43.4 46.3 1202.9 77.6 58.8 73.8 57.0 77.4 + MoCLE 49.3 64.7 46.3 46.9 1222.6 82.1 61.5 78.2 59.8 78.9
+Table 2: Zero-shot results (InstructBLIP) on the held-out datasets, i.e., GQA, VSR, IconQA (IQA), Visdial, MME, POPE and evaluation on held-in datasets, i.e., A-OKVQA, OKVQA, VQAv2.
 Here Direct and MC denote directly answering and multiple choices.
 Best results are marked in bold.
-Models Flickr 30K TextVQA HM SciQA image MSVD QA MSRVTT QA iVQA InstructBLIP (7B) 81.3 53.9 65.3 62.0 41.4 23.0 51.3 81.9 57.1 65.6 63.9 42.6 24.4 53.2 MoCLE (7B) (+0.6) (+3.2) (+0.3) (+1.9) (+1.2) (+1.4) (+1.9)
-Table 2: Zero-shot results on the held-out datasets. Here, HM and SciQA denote HatefulMemes and ScienceQA, respectively.
-4.2 Datasets and Instructions
-To demonstrate the effectiveness of MoCLE, we follow InstructBLIP to use the same training and evaluation data unless mentioned explicitly.
-As Web CapFilt[28], one of the training datasets, is not fully available, we can only download a subset of them.
-Besides this, we utilize all the other training data, including COCO [35], TextCaps [55], VQAv2 [19], OKVQA [45], A-OKVQA [52], OCR-VQA [46], and LLaVA-Instruct150K [38], for instruction tuning.
-For the zero-shot evaluation datasets, we follow InstructBLIP but only keep Flickr30K [64], GQA [22], VSR [36], IconQA [43], TextVQA [56], Hateful Memes [25], ScienceQA [42], MSVD-QA [61], MSRVTT-QA [61] and iVQA [62].
-NoCaps dataset [1] is originally utilized for the image captioning task, which, however, is not adopted here since we cannot reproduce the reported results, probably due to the incomplete Web CapFilt dataset used.
-In addition, we consider another two widely adopted LVLM benchmarks, i.e., MME [16] and POPE [33] for zero-shot evaluations.
-For held-in evaluation, we consider the VQAv2 [19], OKVQA [45] and A-OKVQA [52].
-For ScienceQA, we only evaluate the set with image context.
-Following previous works [14], we report the CIDEr score [57] for Flickr30K, the iVQA accuracy for iVQA, AUC score for HatefulMemes, and Mean Reciprocal Rank (MRR) for Visual Dialog.
-For LVLM benchmarks, we report the perception score for MME and F1 score for the adversarial split of POPE.
+Models Flickr 30K VQAT HM SQA MSVD QA MSRVTT QA iVQA InstructBLIP (7B) 81.3 53.9 65.3 62.0 41.4 23.0 51.3 + MoCLE 81.9 57.1 65.6 63.9 42.6 24.4 53.2
+Table 3: Zero-shot results (InstructBLIP) on the held-out datasets. Here, VQAT, HM and SQA denote TextVQA, HatefulMemes and ScienceQA, respectively.
+Methods Train Data MME MMB SQA GeoQA VQA-RAD SLAKE PathVQA Open Closed Open Closed Open Closed Single LoRA LLaVA-665k 1804 65.89 67.67 - - - - - - - Geo170k - - - 57.82 - - - - - - Med.
+Mix - - - - 53.90 84.19 86.05 85.58 38.07 91.77 Single LoRA All 1794 64.69 66.78 57.56 46.89 77.94 84.61 82.45 35.56 90.71 MoCLE All 1838 66.07 67.38 60.21 53.59 81.98 83.29 85.10 35.21 91.65
+Table 4: Evaluation results of LLaVA-1.5-7B, where MMB denotes MMBench.
+4 Experiment
+In this section, we conduct an assessment of MoCLE across multiple downstream tasks in a zero-shot setting.
+We first detail the experimental settings and implementation details, which are followed by a description of the datasets and instructions employed, along with the outcomes of our evaluations.
+Lastly, we present an ablation study and visualizations of clustering and routing results.
+4.1 Implementation Details
+We evaluate the effectiveness of MoCLE on two LVLMs: InstructBLIP Dai et al.
+(2023) and LLaVA-1.5 Liu et al.
+(2023a).
+Specifically, we compare the performance between the LVMs with and without MoCLE.
+The detailed configuration of MoCLE on these LVLMs are presented in Table 1.
+In addition, we encode all the instructions of different datasets using the all-MiniLM-L6-v2 variant of the Sentence Transformer model Reimers and Gurevych (2019) and cluster their embeddings via $k$-means clustering algorithm.
+More training details can be found in Appendix A.
+4.2 Settings
+For InstructBLIP, we follow Dai et al.
+(2023) for the choice of training datasets.
+However, these datasets only focus on a single domain: natural images.
+To validate the effectiveness of MoCLE on multiple domains, for LLaVA-1.5, in addition to its original training data LLaVA-665K Liu et al.
+(2023a) which focus on natural image domain, we include datasets from multiple domains, i.e., geometric tasks: Geo170k Gao et al.
+(2023a), medical tasks: VQA-RAD Lau et al.
+(2018), SLAKE Liu et al.
+(2021) and PathVQA He et al.
+(2020).
+More details on the training and evaluation datasets are provided in Appendix C.
+Note that during evaluation, we report the CIDEr score Vedantam et al.
+(2015) for Flickr30K, the iVQA accuracy for iVQA, AUC score for HatefulMemes, Mean Reciprocal Rank (MRR) for Visual Dialog, the perception/perception+cognition score for MME (InstructBLIP/LLaVA) and F1 score for the adversarial split of POPE.
 For all other datasets, we report the top-1 accuracy (%).
-The instructions to conduct all the experiments in this paper are from the original InstructBLIP paper.
-The detailed introductions of the held-in and held-out tasks/datasets are also presented in [14].
+Task templates for evaluation can be found in Appendix D.
+LoRA(r=8) r=64 Cluster MoE Uni. Expert LoRA # Params. Flickr HM SQA IQA VSR VQAT Avg.
+(a) 4.19M 81.3 65.1 57.4 44.2 62.8 49.4 60.0
+[…]
+[…]
+[…]
+Table 5: Comparison of individual components of the MoCLE framework in zero-shot vision-language tasks. Default settings are marked in gray.
 4.3 Evaluation Results
-Table 1 and Table 2 present the experimental results on multiple zero-shot and held-in vision-language tasks.
-We build our method upon the Vicuna-7B variant of InstructBLIP.
-Compared to this model, the proposed MoCLE method shows considerable performance improvement.
-Specifically, on held-out datasets such as IconQA, Visual Spatial Reasoning (VSR), TextVQA and ScienceQA datasets, we obtain an absolute performance gain of 2.9% 3.9%, 3.2%, and 1.9%, respectively.
-This indicates that the proposed MoCLE method facilitates generalization to unseen tasks/instructions since these datasets are unseen during training.
-In addition, we observe considerable improvement on common LVLM benchmarks (i.e., MME and POPE).
-POPE evaluates whether the model suffers from object hallucinations while MME considers a wide range of abilities such as math, OCR, color and understanding of arts.
-Therefore, our MoCLE cannot only promote task generation, but also improve overall abilities of the LVLM.
-Further, the held-in tasks whose datasets are seen during training could also benefit from the proposed MoCLE.
-Specifically, with our MoCLE, an absolute improvement of 4.4%, 2.8% and 1.5% could be observed on A-OKVQA (MC), OKVQA and VQAv2.
-This indicates that MoCLE could effectively alleviate the task conflicts during multi-task learning.
+InstructBLIP Tables 2 and 3 show the results on multiple held-out/in vision-language tasks.
+The proposed MoCLE shows considerable performance improvement over the original LVLM.
+Specifically, on held-out datasets such as IconQA, Visual Spatial Reasoning (VSR), TextVQA and ScienceQA datasets, we obtain an absolute performance gain of 2.9%, 3.9%, 3.2%, and 1.9%, respectively.
+On held-in datasets, an absolute improvement of 4.4%, 2.8% and 1.5% can be observed on A-OKVQA (MC), OKVQA and VQAv2, respectively.
+This indicates that the proposed MoCLE facilitates generalization to unseen tasks and can effectively alleviate task conflicts during multi-task learning
+LLaVA-1.5 Similar to the preliminary results in Figure 1, we consider a Single-LoRA baseline where a single set of LoRAs are trained on natural images (LLaVA-665k), geometric (Geo170K), medical (Med.
+Mix) and a mixture of all tasks (All).
+As can be seen, due to task conflicts, the model trained on all tasks shows inferior results compared to those trained on only one task.
+However, MoCLE is able to reduce this gap on medical tasks and even offer better performance on natural image (MME, MMB) and geometric tasks (GeoQA) compared to the model trained on one task.
+This shows that MoCLE is effective with the presence of multiple-domain datasets.
 4.4 Ablation Studies
-In this section, we first ablate the effectiveness of the main components (e.g., task experts and universal expert) in the proposed MoCLE method.
-Then we conduct a thorough analysis to study how the proposed MoCLE method would respond to the change of hyper-parameters (e.g., temperature and the number of clusters and task experts).
-Notice that for the ablations, we use Flickr30K (Flickr), Hateful Memes (HM), ScienceQA (SciQA), IconQA, Visual Spatial Reasoning (VSR) and TextVQA for evaluation.
-Components Flickr HM SciQA IconQA VSR TextVQA Avg.
-InstructBLIP 81.3 65.3 62.0 43.4 60.8 53.9 61.1
-+LoRA (rank=8) 81.3 65.1 57.4 44.2 62.8 49.4 60.0
- (rank=64) 81.5 65.2 62.0 43.9 62.6 49.0 60.7
-+Mixture (Cluster) 81.9 65.4 63.3 46.1 58.9 54.9 61.8
-+Universal Expert 81.9 65.6 63.9 46.3 64.7 57.1 63.3
-Table 3: Comparison of individual components of the MoCLE framework in zero-shot vision-language tasks.
-We include three ablation studies that cumulatively add (i) parameter-efficient finetuning of LLM using LoRA of different ranks (higher ranks are ablated, i.e.
-rank=64, but we use rank=8 for ablations and MoCLE), (ii) replacing the single LoRA with its mixture and route inputs by their cluster, and (iii) training with the universal expert.
-Default settings are marked in gray.
-4.4.1 Effects of Different Components.
-We start from vanilla InstructBLIP with Vicuna-7B and add the main components in the proposed MoCLE method step by step.
-As shown in Table 3, naïvely applying LoRA to InstructBLIP could be problematic.
-Although we observe minor performance gain in IconQA and VSR, the peformance on ScienceQA and TextVQA are much inferior to InstructBLIP.
-We speculate that a single LoRA offers very limited flexibility, thus it performs not so well on many tasks.
-Note that we further conduct an experiment with a rank of 64 (much larger than $r=8$ in the default setting of MoCLE).
-However, the resulting model is still outperformed by InstructBLIP with less trainable parameters.
-This indicates that naïvely increasing the capacity of the network cannot mitigate the task conflicts in the training data.
-However, when we replace a single LoRA with the Mixture of LoRAs and route the input based on the cluster of their instructions, the resulting model outperforms its predecessor by a large margin on SciQA, IconQA, and TextVQA.
-Nevertheless, we notice a sharp performance drop in the VSR task.
-This could be due to the fact that instruction-tuned model generalizes to unseen tasks by training on many instructions, while in our case, each expert sees fewer instructions than the original dense model does.
-As shown in Table 3, this issue could be alleviated by the usage of the universal expert, which is updated by the gradients from all the training data.
-The same result can be observed in TextVQA, which requires not only the VQA ability but also additional optical character recognition (OCR) skills.
-Specifically, in the training stage, we have instructions “Question: {}.
-Short answer:” for all VQA datasets and “OCR tokens: {}.
-Write a description for the photo” for TextCaps [55], while the TextVQA [56] instruction is “OCR tokens: {}, Question: {}.
-Short answer:”, requiring LVLMs to combine knowledge from both tasks.
-4.4.2 Universal Expert vs. Top-2 Experts.
-MoCLE shares similarities with the top-2 variant of MoE.
-Specifically, they both activate two experts each time.
-However, the universal expert differs from top-2 MoE by consistently selecting the universal expert, whose contribution is the complement of the top-1 expert.
-As shown in Table 4, the top-2 MoE model yields inferior results compared to MoCLE.
-Therefore the naïvely activating one more expert cannot address the generalization problem.
-Surprisingly, this variant performs even worse than the MoCLE model without the universal expert reported in Table 3.
-This might be resulted from task conflicts with more parameters shared.
-Flickr HM SciQA IconQA VSR TextVQA Avg.
+In this section, we first ablate the effectiveness of the main components (i.e., Cluster MoE and universal expert) in the proposed MoCLE.
+Then we conduct a thorough analysis to study how the proposed MoCLE responds to changes in hyper-parameters (e.g., temperature and the number of clusters and task experts).
+Notice that we use InstructBLIP for all ablations and we report the evaluation results on Flickr30K (Flickr), Hateful Memes (HM), ScienceQA (SQA), IconQA (IQA), Visual Spatial Reasoning (VSR) and TextVQA (VQAT).
+4.4.1 Effects of Different Components
+We start from MoCLE and remove its key component one-by-one to analyze their effect.
+Universal expert.
+Table 5 shows the ablation results when varying different components of MoCLE.
+By comparing rows (d) and (c), we notice that a sharp performance drop in VSR and TextVQA tasks when the universal expert is removed.
+This is due to that instruction-tuned model generalizes to unseen tasks by training on many instructions, while in our case, each expert sees fewer instructions than the dense model.
+For example, task TextVQA with instruction “OCR tokens: {}, Question: {}.
+Short answer:” needs not only VQA ability but also optical character recognition (OCR) skills, which are learned jointly from VQA data formatted as “Question: {}.
+Short answer:” and TextCaps data formatted as “OCR tokens: {}.
+Write a description for the photo”.
+Thus, universal expert is necessary to maintain generalization ability.
+Cluster MoE.
+Comparing rows (c) and (a), we observe performance drop on SQA, IQA, and VQAT when cluster MoE is not used, which indicates it can alleviate task conflicts within a single set of LoRAs between different tasks.
+LoRA rank.
+As can be seen from rows (c), (b) and (a), naïvely increasing the LoRA ranks from 8 to 64 only leads to a small average performance improvement of 0.7%, and thus cannot address task conflicts.
+Instead, promoting task specialization via clustering achieves notable improvement with fewer additional parameter (×$\times 4$ in Cluster MoE versus ×$\times 8$ when increasing the rank to 64).
+4.4.2 Universal Expert vs. Top-2 Experts
+To ablate the proposed universal expert, we remove it and activate one more existing expert.
+i.e., top-$2$ gating.
+We report their performance in Table 6.
+The top-2 MoE model yields inferior results compared to MoCLE and it even performs worse than the MoCLE variant without the universal expert reported in Table 5.
+This can be explained by the intensified conflicts when task experts are shared via top-$2$ gating because now each expert need to learn common feature with other experts.
+However, as the universal expert is shared all the time especially for this purpose, it frees task experts from this duty and thus alleviates the conflicts.
+Flickr HM SQA IQA VSR VQAT Avg.
 Universal 81.9 65.6 63.9 46.3 64.7 57.1 63.3
@@
 Table 6: Ablation study on the universal expert by comparing with either (i) a universal expert that is activated all the time or (ii) expert with the second largest logit, in addition to the top-1 expert.
-Gating strategy Flickr HM SciQA IconQA VSR TextVQA Avg.
-Token 81.7 65.4 61.9 44.0 49.0 46.6 58.1
-Sentence 82.0 65.1 62.3 45.3 56.6 47.0 59.7
-Dataset 80.3 64.6 63.1 45.9 62.6 55.8 62.1
-Cluster 81.9 65.6 63.9 46.3 64.7 57.1 63.3
-Table 5: Ablation study on routing inputs based on different conditions. Default settings are marked in gray.
-4.4.3 Gating Strategies.
-We compare the proposed cluster-conditioned gating strategy with Token-MoLE, Sentence-MoLE and Dataset-MoLE, where MoLE denotes the mixture of LoRA experts by applying the MoE technique to LoRA, Token-MoLE obtains the routing decision based on the hidden representations of each token, and Sentence-MoLE averages the representations of the instruction tokens while excluding the visual tokens.
-Dataset-MoLE is a special case of MoCLE as it treats each dataset as a cluster while MoCLE leverages $k$-means to obtain clusters among all the datasets.
-Note that the universal expert is adopted for all these MoLEs.
-As observed in Table 5, Token-MoLE and Sentence-MoLE give inferior results on the evaluation tasks.
-We speculate that this is because (1) an expert only learns a subset of the tasks, giving poor generalization and (2) similar tasks are not grouped together by the same expert, resulting in task conflicts within that expert.
-The later can be evidenced by the routing visualization in Figure 4, where samples in the same dataset are routed to multiple experts instead of a dedicate one.
-Further, we observe performance drops when replacing $k$-means clustering with statically treating each dataset as a cluster.
-This results from the fact that Dataset-MoLE is less flexible than MoCLE because it can only assign a dataset to one cluster.
+Gating Flickr HM SQA IQA VSR VQAT Avg.
+Token (LLaVA-MoLE) 81.7 65.4 61.9 44.0 49.0 46.6 58.1
+Chen et al. (2024)
+Sentence (Octavius) 82.0 65.1 62.3 45.3 56.6 47.0 59.7
+Chen et al. (2023g)
+Dataset 80.3 64.6 63.1 45.9 57.6 53.4 60.8
+Jang et al. (2023)
+Cluster 81.9 65.4 63.3 46.1 58.9 54.9 61.8
+Table 7: Ablation study on routing inputs based on different input conditions.
+Figure 3: Ablation study on the number of clusters, experts and gate temperature.
+The x-axes of the first and last figures are log-scaled.
+The y-axes are the average performance of Flickr, HM, SQA, IQA, VSR and VQAT.
+4.4.3 Gating Strategies
+We compare the proposed cluster-conditioned gating strategy with existing MoLE methods in Table 7.
+Note that MoLE denotes the mixture of LoRA experts by applying MoE to LoRA.
+Some of these methods adopt different configurations, e.g.., # experts, expert params.
+and ranks.
+For fair comparison, we follow the first row of Table 1 except that universal expert is not enabled in this experiment.
+Token/Sentence-MoLE.
+The former obtains the routing decision based on the hidden representations of each token (adopted by Chen et al.
+(2024)) and the later on the average representations of the instruction tokens while excluding the visual tokens.
+(adopted by Chen et al.
+(2023f)).
+Both of these methods give inferior results on the evaluation tasks.
+We speculate that this is because (1) a sparse expert learns less data than its dense counterpart, leading to lack of task generalization, (2) similar tasks are not grouped together by the same expert, resulting in task conflicts within that expert, which can be verified by the routing visualization in Figure 6, where samples in the same dataset are routed to multiple experts instead of a dedicated one.
+Dataset-MoLE is a special case of MoCLE as it treats each dataset as a cluster while MoCLE leverages $k$-means to achieve this.
+It closely resembles to the dataset expert proposed in Jang et al.
+(2023) except that we assign clusters for a sentence by its distance to cluster centers while they count, to this sentence, the number of closest reference sentences belonging to each dataset.
+Further, for fair comparison, we only use 4 experts but they allocate an expert for each dataset.
+We observe inferior results compared to the proposed cluster routing.
+This results from the fact that Dataset-MoLE is less flexible as it can only assign a dataset to one cluster.
 However, in practice, we observe multiple tasks in a dataset which should be assigned to different clusters.
-(e.g., llava_150k contains reasoning, conversations and captioning, which are assigned to different clusters/experts as in Figure 2(a) and 3(a)).
-Table 6: Ablation study on the number of clusters $K$.
-# Clusters Flickr HM SciQA IconQA VSR TextVQA
-4 81.3 65.2 63.4 45.3 58.3 55.3
-16 82.0 64.3 61.7 46.3 67.2 56.9
-64 81.9 65.6 63.9 46.3 64.7 57.1
-256 82.5 65.6 63.2 46.0 57.0 56.9
-Temperature $\tau$ Flickr HM SciQA IconQA VSR TextVQA
-0.01 81.6 64.7 61.7 46.3 61.2 55.8
-0.05 81.9 65.6 63.9 46.3 64.7 57.1
-0.1 82.3 65.7 63.7 45.7 63.0 56.7
-0.2 81.9 64.8 62.1 45.2 61.3 54.0
-Table 6: Ablation study on the number of clusters $K$.
-Table 7: Ablation study on the temperature $\tau$ in Eq. (1).
-#Task experts Flickr HM SciQA IconQA VSR TextVQA
-2 81.4 64.4 62.8 45.6 58.1 50.9
-4 81.9 65.6 63.9 46.3 64.7 57.1
-8 81.0 65.3 63.5 45.9 65.0 57.4
-Table 8: Ablation study on the number of task experts.
-4.4.4 Number of Clusters.
-The merit of sentence encoding and clustering lies in reducing the complexity of the instruction, offering robust and semantic-relevant routing process.
-The number of clusters $K$ in this process could measure such complexity.
-A extremely small $K$ would diminish the difference of various tasks, result in many different tasks to be processed by the same expert.
-This could increase the chance of task conflicts.
-As shown in Table 7, when we cluster the inputs into 4 groups, the resulting model performs poorly on the evaluation tasks.
+(e.g., llava_150k contains reasoning, conversations and captioning, which are assigned to different clusters/experts as in Figure 5 and 6(a)).
+4.4.4 Number of Clusters
+The number of clusters $K$ controls the granularity of task specialization.
+A very small $K$ would result in many different tasks to be processed by the same expert, and can increase the chance of task conflicts.
+As shown in Figure 3, when we cluster the inputs into 4 groups, the resulting model performs poorly on the evaluation tasks.
 However, as we increase the number of clusters to 16 and 64, we observe considerable performance gains.
-Nevertheless, too large $K$ introduces unnecessary complexity in the routing process, e.g., a paraphrased instruction gets routed to different experts.
-So we use 64 clusters as the default setting.
-4.4.5 Number of Task Experts.
-As demonstrated in Table 8, more task experts usually provides with stronger capacity.
-Specifically, when only 2 task experts are employed, we observe inferior results on VSR and TextVQA.
+However, a $K$ too large (256) introduces unnecessary complexity to the routing process (e.g., a paraphrased instruction gets routed to different experts).
+So we use 64 clusters by default.
+4.4.5 Temperature
+In the proposed MoCLE, the temperature plays an important role in controlling the contribution of the universal expert.
+Specifically, as shown in Eq.
+(1), $\tau$ controls the sharpness of the gate distribution, while the output of the universal expert is weighted by $1-G_{\text{max}}$.
+Therefore, as $\tau$ decreases, $G_{\text{max}}$ increases, and finally the contribution of the universal expert decreases.
+As shown in Figure 3, the results are consistent with our understandings.
+When $\tau$ is either too small (0.01) or large (0.2) can lead to inferior results.
+The temperatures of $0.05$ and $0.1$ seem to achieve a balance between specialization and generalization of the model.
+In the experiments, we use temperature of 0.05 as default.
+4.4.6 Number of Task Experts
+As demonstrated in Figure 3, more task experts usually provides with stronger capacity.
+Specifically, when only 2 task experts are employed, we observe inferior overall results.
 This model has similar capacity to the single LoRA model in Sec.
@@
 When the number of task experts is increased to 4, the performance gets improved.
-Specifically, on VSR and TextVQA, the 4-expert variant outperforms the 2-expert counterpart by 6.6% and 6.2%, respectively.
-Moreover, we also observe 1.2%, 1.1%, and 0.7% performance gains on HM, SciQA and IconQA.
 When the number of task experts becomes 8, it behaves similarly to the 4-expert case, which indicates that the benefit of increasing capacity converges as we use more task experts.
 Hence, we use 4 task experts as the default setting.
-4.4.6 Temperature.
-In the proposed MoCLE method, the value of temperature plays an important role in controlling the contribution of the universal expert.
-Specifically, as shown in Eq.
-(1), $\tau$ controls the sharpness of the gate distribution, while the output of the universal expert is weighted by $1-G_{\text{max}}$.
-Therefore, as $\tau$ decreases, $G_{\text{max}}$ increases, and finally the contribution of the universal expert decreases.
-As shown in Table 7, the results are consistent with our understandings.
-When a $\tau$ is too small, i.e., 0.01, the model generalizes worse to unseen tasks like SciQA and VSR.
-Similarly, when $\tau$ is too large, i.e., $0.2$, the role of task experts is diminished, leading to task conflicts.
-The temperatures of $0.05$ and $0.1$ seem to achieve a balance between specialization and generalization of the model, and we use 0.05 as the default setting for the temperature.
 4.5 Visualizations
-Figure 3: (a) Clustering assignment of the training datasets when $K=64$.
-The labels on the vertical axis indicate the names of the datasets.
-Within each labeled block, multiple rows represent different subsets, each of which is associated with a unique task template.
-The horizontal axis denotes the cluster index to which these subsets are assigned.
-(b) t-SNE visualization of the instruction encoding.
-Different colors correspond to different datasets, while the shape of the markers indicates the task category defined manually.
-4.5.1 Clustering.
+Figure 4: T-SNE visualization of the instruction encoding. Different colors correspond to different datasets, while the shape of the markers indicates the task category defined manually.
+Figure 5: Clustering assignment of the training datasets when $K=64$.
+The labels on the y-axis indicate the names of the datasets.
+The x-axis denotes the cluster index to which the subsets are assigned.
+4.5.1 Clustering
 We first show the justification to represent the training data via their instructions.
-Specifically, for each dataset, we sample 100 examples and encode their instructions with the all-MiniLM-L6-v2 variant of the Sentence Transformer model [49].
-After that, we visualize the data in Figure 2(b) via t-SNE [44].
-According to Figure 2(b), we observe that (1) samples from different datasets but from the same task are grouped together.
-For example, coco_vqg, aok_vqg, and ok_vqg are all VQG (visual question generation) tasks (triangle markers) and reside on the left part of the figure.
-(2) samples from similar tasks are closed to each other.
-For example, coco_cap and textcaps both belong to the image captioning (CAP) task (small dots) and stay close to each other.
-Similarly, both the visual question answering (VQA, cross markers) task with datasets (e.g., aok_vqa, coco_vqa, ocrvqa, and ok_vqa) and conversation (CONV) task with datasets (e.g., llava_conversation, llava_detail, and llava_reason) involve answering user questions, which lie in the middle part of Figure 2(b), suggesting that instructions are good representatives of the training data.
+Specifically, for each dataset, we sample 100 examples and encode their instructions with the all-MiniLM-L6-v2 variant of the Sentence Transformer model Reimers and Gurevych (2019).
+We then visualize the data in Figure 4 via t-SNE Van der Maaten and Hinton (2008).
+As can be seen, (1) Samples from the same task are grouped together.
+For example, all visual question generation (VQG, triangle markers) data reside on the left part of the figure.
+(2) Samples from similar tasks are close to each other, e.g., coco_cap and textcaps both belong to the image captioning (CAP, small dots) task and stay close to each other at the lower right of the figure.
+Similarly, both visual question answering (VQA, cross markers) and conversation (CONV, y-shape) data involve answering user questions, which lie in the middle part of the figure, suggesting that instructions are good representatives of training data.
 We then cluster all the instructions of the examples in the training data into 64 groups using $k$-means clustering.
-Figure 2(a) shows the cluster assignment of the training data.
+Figure 5 shows the cluster assignment of the training data.
 Here, each row in the heatmap denotes a subset of a dataset.
@@
 3.1) on the samples of the dataset.
-We observe (1) different subsets of the same datasets are assigned to similar clusters.
-For example, aok_vqa, coco_vqa, and ok_vqa are in the first several clusters, while aok_vqg, coco_vqg, and ok_vqg mainly occupy 4 other clusters.
+We observe the following: (1) Different subsets of the same datasets are assigned to similar clusters.
+For example, aok_vqa, coco_vqa, and ok_vqa are in the first several clusters.
 (2) Datasets of similar tasks are assigned to common clusters.
-For example, llava_150k including llava_detail, llava_detail and llava_conversation and a series of VQA tasks share the first several clusters as they are to answer questions.
-Those evidences justify the use of clustering on task instructions as an automatic partition strategy for training datasets.
-4.5.2 Routing results.
-Figure 4 visualizes the routing decisions of our MoCLE and Sentence-MoLE.
+For example, llava_150k including llava_detail, llava_reason and llava_conversation and a series of VQA tasks share the first several clusters as they are to answer questions.
+These justify the use of clustering on task instructions as an automatic partition strategy for training datasets.
+4.5.2 Routing Results
+Figure 6 visualizes the routing decisions of the proposed MoCLE and Sentence-MoLE.
 We obtain both results from one mixture of LoRA, i.e., one linear module in a layer.
-The routing results are aggregated by the subset of datasets similar to Figure 2(a).
-Shown in Figure 3(a), MoCLE could achieve task-level routing for the inputs.
-For example, datasets from VQA and VQG tasks are handled by expert 0 and 3 respectively.
-In addition, VQA expert (i.e., expert 0, which is mostly occupied by VQA task) is fundamental for most of the datasets.
-That is, most held-out datasets activate the VQA expert.
-This could be justified by the fact that most vision-language tasks take the form of question answering.
-However, this does not indicate that the entire model collapses to a single expert because knowledge from other tasks are learned by the universal expert and are leveraged during evaluation.
-Further, the remaining 3 experts are not dispensable as without them all tasks would clutter in a single expert, resulting in task conflicts.
-Instead, routing pattern of Sentence-MoLE in Figure 3(b) reveals little correlations between tasks/datasets and experts.
-That is, different datasets/tasks obtain similar routing decisions, still suffering from task conflicts.
+The routing results are aggregated by the subset of datasets similar to Figure 5.
+As can be seen from Figure 6(a), MoCLE can achieve task-level routing for the inputs.
+For example, datasets from VQA and VQG tasks are handled by expert 0 and 3, respectively.
+Instead, routing pattern of Sentence-MoLE in Figure 6(b) reveals little correlations between datasets and experts.
+That is, different datasets obtain similar routing decisions, and thus still suffer from task conflicts.
 (a) Our MoCLE.
@@
 Figure 6: Routing decisions of one LoRA mixture for MoCLE and Sentence-MoLE.
-The setup of the vertical axis is similar to Figure 2(a) except that we also include the held-out tasks.
+The setup of the vertical axis is similar to Figure 5 except that we also include the held-out tasks.
 They are separated by a dotted line on the vertical axis.
@@
 5 Conclusions
-In this paper, we first show that the task conflicts exists in vision language instruction tuning through extensive experiments.
+In this paper, we first show through extensive experiments that task conflicts exist in vision language instruction tuning.
 To address this, we propose the Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel MoE architecture designed to activate the task-customized model parameters based on the instruction clusters.
 In addition, we achieve task specialization and generalization in MoCLE simultaneously via a separate universal expert.
-Comprehensive evaluations of MoCLE on 11 unseen tasks show the effectiveness of our MoCLE.
-Acknowledgement
+Comprehensive evaluations of MoCLE on both held-out/in tasks show the effectiveness of MoCLE.
+6 Limitations
+Although effective, we mainly focus on task conflicts among text-based conversation tasks in this paper, while the support of our MoCLE for more complicated visual perception tasks is appealing, which has shown more severe task conflicts with the conversation tasks Zhu et al.
+(2022).
+7 Acknowledgement
 We gratefully acknowledge the support of MindSpore, CANN (Compute Architecture for Neural Networks) and Ascend AI Processor used for this research.
-Appendix
-Appendix 0.A Sources of Results in Table 1
-For LLaVA (13B) and MiniGPT-4 (13B), the results on GQA, VSR, OKVQA and IconQA are from the Table 3 of [5], the results on POPE and A-OKVQA (Direct) are from Table 3 and 6 of [33], respectively.
-In particular, the performance of MiniGPT-4 (13B) on MME is from its official leaderboard  [^1] https://github.com/BradyFU/Awesome-Multimodal-Large-Language-Models/tree/Evaluation.
-All the task performances of InstructBLIP (7B) are obtained from its reproduced version implemented in this paper.
-Appendix 0.B Training Details
-Following [14], we adopt the same training configurations for the mentioned models such as the proposed MoCLE, the reproduced InstructBLIP (7B) and the task experts in Sec.
+Appendix A Training Details
+A.1 InstructBLIP
+Following Dai et al.
+(2023), we adopt the same training configurations for the mentioned models such as the proposed MoCLE, the reproduced InstructBLIP (7B) and the task experts in Sec.
 1.
 We train those models with a maximum of 60K steps and a batch size of 128.
-The AdamW optimizer [26] is used, with $\beta_{1}$ as 0.9, $\beta_{2}$ as 0.999, and a weight decay as 0.05.
+The AdamW optimizer Kingma and Ba (2014) is used, with $\beta_{1}$ as 0.9, $\beta_{2}$ as 0.999, and a weight decay as 0.05.
 We apply a linear warmup of the learning rate during the initial 1000 steps, increasing from ${10}^{-8}$ to ${10}^{-5}$, followed by a cosine decay with a minimum learning rate of 0.
-Appendix 0.C Weights of the Universal Experts
+A.2 LLaVA-1.5
+We follow Liu et al.
+(2023a) for the training configuration.
+Specifically, for LLaVA-150KLiu et al.
+(2023a), Geo170KGao et al.
+(2023a) and Med.
+Mix, i.e., VQA-RADLau et al.
+(2018), SLAKELiu et al.
+(2021) and Path-VQAHe et al.
+(2020), we train the model for 1, 2 and 9 epochs, respectively.
+When training on all of these datasets, we copy each dataset $k$ times ($k$ is the number of epochs it is trained independently) and merge them into a single dataset.
+For each training job, we use a batch size of 128, weight decay of 0 and learning rate of $41e-$ 4, which is warmed up from 0 during the initial 3% steps and followed by a cosine decay with a minimum learning rate of 0.
+Appendix B Weights of the Universal Experts
 During training, if some training data obtains a very large weight on a task expert, such data tend to be very specific and might be less beneficial to other tasks.
@@
 Figure 7: Weights of the universal expert for different datasets. Colors indicate different datasets.
-Appendix 0.D Datasets for Vision-language Instruction Tuning
-Here we give a detailed description of datasets used during instruction tuning.
+Appendix C Data
+We list the training and evaluation data for InstructBLIP and LLaVA-1.5 in Table 8.
+C.1 Dataset Abbreviation
+For InstructBLIP, we further show in Table 9 (1) the abbreviation used in Sec.
+4.5 for each dataset and (2) their manually defined task category.
 As shown in Table 9, we use 13 datasets in total.
-Here multiple datasets might be associated with the same data sources because these sources are formatted by different groups of task templates (see Appendix 0.E).
-For LLaVA-Instruct150K [38], we do not apply any task template as it has been well formatted.
-For Web Capfilt, we use COCO+VG+CC+SBU (14M images) as [29].
-Regarding with which split is used, please refer to the InstructBLIP paper [14] for more details.
-During training, we sample datasets using the data balancing strategy as in [14].
-Appendix 0.E Task Templates
-Following [14], we use the same set of task templates for instruction tuning and held-in/out evaluation. Please refer to Tables 14 and 15 for training and evaluation templates.
-Appendix 0.F Case Studies
+Here multiple datasets might be associated with the same data sources because these sources are formatted by different groups of task templates (see Appendix D).
+For LLaVA-150K Liu et al.
+(2023b), we do not apply any task template as it has been well formatted.
+Models Training datasets Evaluation Datasets InstructBLIP Flickr30K Young et al.
+(2014), GQA Hudson and Manning (2019) Web CapFilt Li et al.
+(2023a) VSR Liu et al.
+(2022a), IconQA Lu et al.
+(2021) A-OKVQA Schwenk et al.
+(2022) TextVQA Singh et al.
+(2019), Hateful Memes Kiela et al.
+(2020) TextCaps Sidorov et al.
+(2020), VQAv2Goyal et al.
+(2017) ScienceQA Lu et al.
+(2022), MSVD-QA Xu et al.
+(2017) OKVQA Marino et al.
+(2019), COCO Lin et al.
+(2014) MSRVTT-QA Xu et al.
+(2017), iVQA Yang et al.
+(2021) OCRVQA Mishra et al.
+(2019), LLaVA-150K Liu et al.
+(2023b) MME Fu et al.
+(2023), POPE Li et al.
+(2023d) OKVQA, A-OKVQA, VQAv2 LLaVA-1.5 LLaVA-665K Liu et al.
+(2023a) MME, MMBenchLiu et al.
+(2023c), ScienceQA Geo170K Gao et al.
+(2023a), VQA-RAD Lau et al.
+(2018) GeoQA Chen et al.
+(2021a), VQA-RAD SLAKE Liu et al.
+(2021), PathVQA He et al.
+(2020) SLAKE, PathVQA
+Table 8: Datasets used for training and evaluation. *: the train split of this dataset is used during instruction tuning.
+Datasets Data Source Task Template Group aok_vqa A-OKVQA VQAMC aok_vqg A-OKVQA VQG coco_cap COCO CAP coco_vqa VQAv2 VQA coco_vqg VQAv2 VQG ocrvqa OCR-VQA VQA ok_vqa OKVQA VQA ok_vqg OKVQA VQA textcaps TextCaps OCRCAPS capfilt Web CapFilt CAP llava_conversation LLaVA-150K - llava_detail LLaVA-150K - llava_reason LLaVA-150K -
+Table 9: Abbreviation and manually defined task categories for the training datasets of InstructBLIP.
+Appendix D Task Templates
+For InstructBLIP, we use the same set of task templates following Dai et al.
+(2023) for instruction tuning and held-in/out evaluation.
+Please refer to Tables 10 and 11 for training and evaluation templates
+Template Group Task Template CAP A short image caption: A short image description: A photo of An image that shows Write a short description for the image.
+Write a description for the photo.
+Provide a description of what is presented in the photo.
+Briefly describe the content of the image.
+Can you briefly explain what you see in the image?
+Could you use a few words to describe what you perceive in the photo?
+Please provide a short depiction of the picture.
+Using language, provide a short account of the image.
+Use a few words to illustrate what is happening in the picture.
+VQA {Question} Question: {Question} Question A short answer to the question is Q: {Question} A: Question: {Question} Short answer: Given the image, answer the following question with no more than three words.
+{Question} Based on the image, respond to this question with a short answer: {Question}.
+Answer: Use the provided image to answer the question: {Question} Provide your answer as short as possible: What is the answer to the following question?
+“{Question}” The question “{Question}” can be answered using the image.
+A short answer is VQG Given the image, generate a question whose answer is: {Answer}.
+Question: Based on the image, provide a question with the answer: {Answer}.
+Question: Given the visual representation, create a question for which the answer is “{Answer}”.
+From the image provided, craft a question that leads to the reply: {Answer}.
+Question: Considering the picture, come up with a question where the answer is: {Answer}.
+Taking the image into account, generate an question that has the answer: {Answer}.
+Question:
+[…]
+Evaluation Dataset Instruction
+GQA, iVQA, Question: {} Short answer:
+MSVD, MSRVTT,
+OKVQA, VQAV2,
+A-OKVQA(Direct)
+A-OKVQA(MC), Question: {} Options: {}. Short answer:
+IconQA
+NoCaps, Flickr30k A short image description:
+TextVQA OCR tokens: {}. Question: {} Short answer:
+ScienceQA Context: {} Question: {} Options: {}. Answer:
+HatefulMemes This is an image with: “{}” written on it. Is it hateful? Answer:
+VSR Based on the image, is this statement true or false? “{}” Answer:
+[…]
+Table 11: Task templates used during evaluation.
+Appendix E Case Studies
 In this section, we present several case studies with MoCLE.
@@
 Then we showcase some example instructions sampled from different clusters.
-0.F.1 Conversations
+E.1 Conversations
 In Table 12, we instruct the model to conduct a very difficult object counting task.
 The correct answer for this question is 63, which is quite hard for existing LVLMs.
-InstructBLIP fails to give the correct answers, while with MoCLE, InstructBLIP can respond the user query in a much more proper manner.
+InstructBLIP fails to give the correct answers, while with MoCLE, InstructBLIP can respond the user query in a more proper manner.
 In Table 13, the model is queried to recognize the character in the image.
@@
 Similarly, in Table 15, we ask the model to describe the image in a detailed manner.
-InstructBLIP still mistakes this query as an image caption task and gives very short caption to this image.
-Instead, with MoCLE, the model correctly understands the “in details” in the instruction and provides sufficient details.
-Datasets Data Source Task Template Group
-aok_vqa A-OKVQA [52] VQAMC
-aok_vqg A-OKVQA [52] VQG
-coco_cap COCO [35] CAP
-coco_vqa VQAv2 [19] VQA
-coco_vqg VQAv2 [19] VQG
-ocrvqa OCR-VQA [46] VQA
-ok_vqa OKVQA [45] VQA
-ok_vqg OKVQA [45] VQA
-textcaps TextCaps [55] OCRCAPS
-capfilt Web CapFilt [29] CAP
-llava_conversation LLaVA-Instruct150K [38] -
-llava_detail LLaVA-Instruct150K [38] -
-llava_reason LLaVA-Instruct150K [38] -
-Table 9: List of datasets used for training.
+InstructBLIP still mistakes this query as an image caption task and gives a very short caption.
+Instead, with MoCLE, the model correctly understands the “in details” in the instruction and gives sufficient details.
 Model responses to counting-related queries.
-[Uncaptioned image] User How many books are there?
-New Bing I see that you have sent me an image.
-According to the description you provided, there are 30 books on the bookshelf.
-InstructBLIP 100 +MoCLE There are a lot of books on the shelves.
-It is difficult to say exactly how many books there are, but it appears to be a large number.
+User How many books are there?
+New Bing I see that you have sent me an image. According to the description you provided, there are 30 books on the bookshelf.
+InstructBLIP 100
++MoCLE There are a lot of books on the shelves. It is difficult to say exactly how many books there are, but it appears to be a large number.
 Table 12: Model responses to counting-related queries.
 Model responses to OCR-related queries.
-[Uncaptioned image] User What are all the exact scene texts in the image?
+User What are all the exact scene texts in the image?
 New Bing According to the description you provided, the image contains a license plate with the text “S3302CD”.
 If you have any other questions or if there is anything else I can help you with, please let me know!
-InstructBLIP The exact characters in the image are “S 3000 CD.” +MoCLE The exact characters in the image are “S 3302 CD.” Table 11: Model responses to OCR-related queries.
+InstructBLIP The exact characters in the image are “S 3000 CD.”
++MoCLE The exact characters in the image are “S 3302 CD.”
+Table 13: Model responses to OCR-related queries.
 Model responses when asked to introduce a famous person.
-[Uncaptioned image] User Can you give an introduction to this person?
-New Bing I’m sorry, I cannot give an introduction to this person as I am not allowed to describe people’s identity, gender, race or emotions InstructBLIP albert einstein +MoCLE Albert Einstein was a German-born theoretical physicist who developed the theory of relativity.
-He is considered one of the most influential scientists of the 20th century.
+User Can you give an introduction to this person?
+New Bing I’m sorry, I cannot give an introduction to this person as I am not allowed to describe people’s identity, gender, race or emotions
+InstructBLIP albert einstein
++MoCLE Albert Einstein was a German-born theoretical physicist who developed the theory of relativity. He is considered one of the most influential scientists of the 20th century.
 Table 14: Model responses to introduce a celebrity.
-0.F.2 Sample Instructions in Clusters
+E.2 Sample Instructions in Clusters
 In Table 16, we showcase some sample instructions assigned to different clusters.
 Though all the instructions in the 4 selected clusters belong to VQA-related tasks, they focus on various perspectives such as food, pet, men, and counting, justifying the usage of a large number of instruction clusters.
-Moreover, in Table 7, we show that the model performs poorly with only a few clusters, i.e., $K=4$, and gets better as we increase $K$.
-With a few clusters, VQA, VQG, CONV and CAP tasks will occupy 4 clusters, leaving no more space for samples with different topics.
-Thus, the model might perform worse when mixing these topics.
+Appendix F Future Works
+Although effective, we mainly focus on task conflicts among text-based conversation tasks (e.g., image captioning and VQA Gou et al.
+(2023)) in this paper, while the support of our MoCLE for more complicated visual perception tasks (e.g., object detection Han et al.
+(2021); Li et al.
+(2022)) is appealing, which has shown more severe task conflicts with conversations Zhu et al.
+(2022).
+Moreover, the applicability of MoCLE to visual pre-training Chen et al.
+(2021b, 2023b) and diffusion models Chen et al.
+(2023d); Gao et al.
+(2023b); Liu et al.
+(2023d); Li et al.
+(2023c) is also a potential research direction, considering the ubiquity of negative transfer phenomenon.
 Model responses when queried to give detailed image descriptions.
-[Uncaptioned image] User Introduce me this image in details.
+User Introduce me this image in details.
 New Bing It appears to be a photograph of an adult elephant walking on a beach.
@@
 The elephant is walking on the sand and there are a few birds in the foreground.
-The elephant appears to be dark grey in color and has large ears and tusks InstructBLIP an elephant walking along the beach +MoCLE An elephant is walking along the beach, with birds in the background.
-The elephant’s trunk is extended towards the water, and it appears to be enjoying its time at the beach.
+The elephant appears to be dark grey in color and has large ears and tusks
+InstructBLIP an elephant walking along the beach
++MoCLE An elephant is walking along the beach, with birds in the background. The elephant’s trunk is extended towards the water, and it appears to be enjoying its time at the beach.
 Table 15: Model responses to give detailed image description.
-Template Group
-Task Template
-CAP
-“A short image caption:”
-“A short image description:”
-“A photo of”
-“An image that shows”
-“Write a short description for the image.”
-“Write a description for the photo.”
-“Provide a description of what is presented in the photo.”
-“Briefly describe the content of the image.”
-“Can you briefly explain what you see in the image?”
-“Could you use a few words to describe what you perceive in the photo?”
-“Please provide a short depiction of the picture.”
-“Using language, provide a short account of the image.”
-“Use a few words to illustrate what is happening in the picture.”
-VQA
-“{Question}”
-“Question: {Question}”
-“Question A short answer to the question is”
-“Q: {Question} A:”
-“Question: {Question} Short answer:”
-“Given the image, answer the following question with no more than three words. {Question}”
-“Based on the image, respond to this question with a short answer: {Question}. Answer:”
-“Use the provided image to answer the question: {Question} Provide your answer as short as possible:”
-“What is the answer to the following question? {Question}”
-“The question {Question} can be answered using the image. A short answer is”
-VQG
-“Given the image, generate a question whose answer is: {Answer}. Question:”
-“Based on the image, provide a question with the answer: {Answer}. Question:”
-“Given the visual representation, create a question for which the answer is {Answer}.”
-“From the image provided, craft a question that leads to the reply: {Answer}. Question:”
-“Considering the picture, come up with a question where the answer is: {Answer}.”
-“Taking the image into account, generate an question that has the answer: {Answer}. Question:”
-[…]
-Evaluation Dataset
-Instruction
-GQA, iVQA,
-“Question: {} Short answer:”
-MSVD, MSRVTT,
-OKVQA, VQAV2,
-A-OKVQA(Direct)
-A-OKVQA(MC),
-“Question: {} Options: {}. Short answer:”
-IconQA
-NoCaps, Flickr30k
-“A short image description:”
-TextVQA
-“OCR tokens: {}. Question: {} Short answer:”
-ScienceQA
-“Context: {} Question: {} Options: {}. Answer:”
-HatefulMemes
-“This is an image with: “{}” written on it. Is it hateful? Answer:”
-VSR
-“Based on the image, is this statement true or false? “{}” Answer:”
-Visdual Dialog
-[…]
-Table 15: Task templates used during evaluation.
-Cluster
-Instruction Samples
-Topics
-1
-“Q: what is being done to the food in the glass fronted box? A:”
-Food
+Cluster Instruction Samples Topics
+1 “Q: what is being done to the food in the glass fronted box? A:” Food
 “Q: what category of pizza would this fall into? Options: (a) vegetarian (b) meat lovers (c) pesto (d) pepperoni, A:”
@@
 “what category of pizzas would this be considered?”
-2
-“Q: what sport is the cartoon dog playing? A:”
-Pet
+2 “Q: what sport is the cartoon dog playing? A:” Pet
 “Question: what is likely her favorite animal? Options: (a) cat (b) dog (c) pig (d) sheep, Short answer:”
@@
 “What might the relationship between the two women and the dog be?”
-3
-“what type dressing does this man favor?”
-Men
+3 “what type dressing does this man favor?” Men
 “Based on the image, respond to this question with a short answer: what are the men doing?. Answer:”
@@
 “Question: why is the man kneeling on the ground?”
-4
-“how many more animals need to be added to all of these to get the number ten?”
-Counting
+4 “how many more animals need to be added to all of these to get the number ten?” Counting
 “Question: how many big elephants are inside of this zoo enclosure together? Options: (a) one (b) four (c) two (d) three, Short answer:”
@@
 Table 16: Sampled instructions from different clusters.
-Appendix 0.G Limitation
-Although effective, we mainly focus on task conflicts among text-based conversation tasks (e.g., image captioning and VQA [18]) in this paper, while the support of our MoCLE for more complicated visual perception tasks (e.g., object detection [20, 30]) is appealing, which has shown more severe task conflicts with the conversation tasks [74].
-Moreover, the applicability of MoCLE to image diffusion models [9, 17, 39, 32] is also a potential research direction, considering the ubiquity of negative transfer phenomenon.